inputenc.sty

Alan Jeffrey  Frank Mittelbach

v1.0d 2004/02/05

1 Introduction

This package allows the user to specify an input encoding (for example, ASCII, ISO Latin-1 or Macintosh) by saying:

  \usepackage[encoding name]{inputenc}

The encoding can also be selected in the document with:

  \inputencoding{encoding name}

Originally this command was only to be used in vertical mode (with the idea that it should be only within a document when using text from several documents to build up a composite work such as a volume of journal articles. However, usages in certain languages suggested that it might be preferable to allow changing the input encoding at any time, which is what is possible now (though that is quite computing resource intensive).

The encodings provided by this package are:

• ascii ASCII encoding for the range 32–127.

• latin1 ISO Latin-1 encoding.

• latin2 ISO Latin-2 encoding.

• latin3 ISO Latin-3 encoding.

• latin4 ISO Latin-4 encoding.

• latin5 ISO Latin-5 encoding.

• latin9 ISO Latin-9 encoding.

• latin10 ISO Latin-10 encoding.

• decmulti DEC Multinational Character Set encoding.

• cp850 IBM 850 code page.

• cp852 IBM 852 code page.

• cp858 IBM 858 code page (this is 850 with euro symbol).

• cp437 IBM 437 code page.

• cp437de IBM 437 code page (German version).

• cp865 IBM 865 code page.

• applemac Macintosh encoding.

• macce Macintosh Central European codepage.

• next Next encoding.

• cp1250 Windows 1250 (central and eastern Europe) code page.

• cp1252 Windows 1252 (Western Europe) code page.

• cp1257 Windows 1257 (Baltic) code page.

• ansinew Windows 3.1 ANSI encoding, extension of Latin-1 (synonym^{1} for

     cp1252).

Each encoding has an associated .def file, for example latin1.def which defines the behaviour of each input character, using the commands:

     \DeclareInputText{slot}{text}
     \DeclareInputMath{slot}{math}

This defines the input character slot to be the text material or math material respectively. For example, latin1.def defines slots "D6 (Æ) and "B5 (μ) by saying:

    \DeclareInputText{214}{\AE}
    \DeclareInputMath{181}{\mu}

Note that the commands should be robust, and should not be dependent on the output encoding. The same slot should not have both a text and a math declaration for it. (This restriction may be removed in future releases of inputenc).

The .def file may also define commands using the declarations:

\providecommand or \ProvideTextCommandDefault. For example:
    \ProvideTextCommandDefault{\textonequarter}{\ensuremath{\frac14}}
    \DeclareInputText{188}{\textonequarter}

The use of the ‘provide’ forms here will ensure that a better definition will not be over-written; their use is recommended since, in general, the best definition depends on the fonts available.

See the documentation in fntguide.tex and ltoutenc.dtx for details of how to declare text commands.

1It is now generated using the guards cp1252,ansinew the latter only used for the provides file line.

2 Announcing the files

We announce the files:

 1 ⟨package⟩\NeedsTeXFormat{LaTeX2e}[1995/12/01]
 2 ⟨package⟩\ProvidesPackage{inputenc}
 3 ⟨ascii⟩ \ProvidesFile{ascii.def}
 4 ⟨latin1⟩ \ProvidesFile{latin1.def}
 5 ⟨latin2⟩ \ProvidesFile{latin2.def}
 6 ⟨latin3⟩ \ProvidesFile{latin3.def}
 7 ⟨latin4⟩ \ProvidesFile{latin4.def}
 8 ⟨latin5⟩ \ProvidesFile{latin5.def}
 9 ⟨latin9⟩ \ProvidesFile{latin9.def}
10 ⟨latin10⟩ \ProvidesFile{latin10.def}
11 ⟨decmulti⟩ \ProvidesFile{decmulti.def}
12 ⟨cp850⟩ \ProvidesFile{cp850.def}
13 ⟨cp852⟩ \ProvidesFile{cp852.def}
14 ⟨cp858⟩ \ProvidesFile{cp858.def}
15 ⟨cp437⟩ \ProvidesFile{cp437.def}
16 ⟨cp437de⟩ \ProvidesFile{cp437de.def}
17 ⟨cp865⟩ \ProvidesFile{cp865.def}
18 ⟨applemac⟩ \ProvidesFile{applemac.def}
19 ⟨applemacce⟩ \ProvidesFile{macce.def}
20 ⟨next⟩ \ProvidesFile{next.def}
21 ⟨ansinew⟩ \ProvidesFile{ansinew.def}
22 ⟨cp1252&!ansinew⟩ \ProvidesFile{cp1252.def}
23 ⟨cp1250⟩ \ProvidesFile{cp1250.def}
24 [2004/02/05 v1.0d Input encoding file]

25 ⟨cp850⟩%%

26 ⟨cp850⟩%% If you need a euro symbol, try cp858 instead.

27 ⟨cp850⟩%%

Now we insert a \makeatletter at the beginning of all .def files.

28 ⟨−package⟩\makeatletter

3 The package

Before we start with the code, an important comment is in order: as you may or may not know, the tabbing environment changes the definition of the commands \', \`, and \=. Outside such an environment these commands produce the corresponding accents, inside they are used for special text positioning, and the accents can be accessed using \a', \a`, and \a=. Therefore we must use the latter instead of the former in the second argument to \DeclareInputText, e.g.

(from latin1.def):
     \DeclareInputText{224}{\@tabacckludge`a}

The command \@tabacckludge is defined (in ltoutenc.dtx) in such a way that \@tabacckludge' will expand to the internal form of \'. Thus it is \' that is carried around internally (the same applies to the other two accent commands).

\DeclareInputText \DeclareInputMath \IeC These commands declare the expansion of an active character. The math declaration is the usual trick with \uppercase. The text declaration is sneakier, since in text space matters. We look to see if the definition ends in a macro, by checking whether it’s \meaning ends in a space. If it does, then we add an irrelevant \IeC and braces around the definition, in order to avoid any space after the active char being gobbled up once the text is written out to an auxiliary file.

The definition should contain only robust commands (and, for correct ligatures and kerning, they must be defined via the interfaces in the fontenc package).

29 ⟨∗package⟩

30 \def\DeclareInputMath#1{%
31 \@inpenc@test
32 \bgroup
33  \uccode`\~#1%
34  \uppercase{%
35 \egroup
36  \def~%
37 }%

38 }

39 \def\DeclareInputText#1#2{%
40 \def\reserved@a##1 ${}%
41 \def\reserved@b{#2}%
42 \ifcat_\expandafter\reserved@a\meaning\reserved@b$ $_%
43  \DeclareInputMath{#1}{#2}%
44 \else
45  \DeclareInputMath{#1}{\IeC{#2}}%
46 \fi

47 }

The definition of \IeC was modified not to insert a \protect unless it is needed, this means it works in \hyphenation commands, and other such delicate places. It was then further changed to never insert a \protect as one is never needed; this makes it work in even more places.

This still needs some attention.

48 \def\IeC{%
49 \ifx\protect\@typeset@protect
50  \expandafter\@firstofone
51 \else
52  \noexpand\IeC
53 \fi

54 }

\inputencoding This sets the encoding to be #1. It first sets all the characters 128–255 to be active (and sets their initial definition to be \@inpenc@undefined). It now also does this for some ‘low’ codes below 32, but misses out Null, control-I, control-J, control-L and control-M.

It then inputs #1.def. But it first sets up a test that produces a warning message if no suitable definitions get read.

55 \def\inputencoding#1{%
56 \gdef\@inpenc@test{\global\let\@inpenc@test\relax}%

Keyboard characters which don’t get a definition will be mapped to \@inpenc@undefined which gets a definition producing an error message indicating in which input encoding the current keyboard character is undefined:

57 \edef\@inpenc@undefined{\noexpand\@inpenc@undefined@{#1}}%

The \edef in the above definition is essential as #1 may be \CurrentOption in which case a later use would return incorrect information (at best nothing).

Now we make all potential input characters active.

58 \@inpenc@loop\^^A\^^H%
59 \@inpenc@loop\^^K\^^K%
60 \@inpenc@loop\^^N\^^_%
61 \@inpenc@loop\^^?\^^ff%

To be able to process the input encoding file in horizontal mode we need to ensure that we don’t get any stray spaces into the horizontal mode or else we end up with extra space in the paragraph.

62 \advance\endlinechar\@M
63 \xdef\saved@space@catcode{\the\catcode`\ }%
64 \input{#1.def}%
65 \advance\endlinechar-\@M
66 \catcode`\ \saved@space@catcode\relax
   If there have been no \DeclareInputText or \DeclareInputMath commands

read then something is amiss.

67 \ifx\@inpenc@test\relax\else
68  \PackageWarning{inputenc}%
69  {No characters defined\MessageBreak
70  by input encoding change to `#1'\MessageBreak}%
71 \fi

72 }

\@inpenc@undefined@ This command will assigned to any active character unless it get a proper definition by the encoding. The argument is the current encoding name.

73 \def\@inpenc@undefined@#1{\PackageError{inputenc}%
74  {Keyboard character used is undefined\MessageBreak
75  in inputencoding `#1'}%
76  {You need to provide a  definition with
77  \noexpand\DeclareInputText\MessageBreak or
78  \noexpand\DeclareInputMath before using this key.}}%

\@inpenc@loop Make characters #1 to #2 inclusive active and undefined.

79 \def\@inpenc@loop#1#2{%
80 \@tempcnta`#1\relax
81 \loop
82  \catcode\@tempcnta\active
83  \bgroup
84  \uccode`\~\@tempcnta
85  \uppercase{%
86  \egroup
87  \let~\@inpenc@undefined
88  }%
89 \ifnum\@tempcnta<`#2\relax
90  \advance\@tempcnta\@ne
91 \repeat}

Then for each option, we input that encoding file.

92 \DeclareOption*{\inputencoding{\CurrentOption}}
93 \ProcessOptions

94 ⟨/package⟩

4 Default definitions for characters

Some input characters map to internal functions which are not in either the T1 or OT1 font encoding. For this reason default definitions are provided in the encoding file: these will be used unless some other output encoding is used which supports those glyphs. In some cases this default definition has to be simply an error message.

Note that this works reasonably well only because the encoding files for both OT1 and T1 are loaded in the standard LaTeX format.

The name \textblacksquare is derived from the AMS symbol name since Adobe seem not to want this symbol. The default definition, as a rule, makes no claim to being a good design.

Some entries are repeated in case guards must all be on one line.

95 ⟨∗latin1 | decmulti | latin2 | latin3 | latin4 | latin5 | latin9 | applemacce | latin10⟩

 96 \ProvideTextCommandDefault{\textdegree}{\ensuremath{{^\circ}}}

97 ⟨/latin1 | decmulti | latin2 | latin3 | latin4 | latin5 | latin9 | applemacce | latin10⟩ 98 ⟨∗cp850 | cp858 | cp852 | cp865 | cp437 | cp437de | applemac | cp1252 | cp1250 | cp1257 | next⟩

 99 \ProvideTextCommandDefault{\textdegree}{\ensuremath{{^\circ}}}

100 ⟨/cp850 | cp858 | cp852 | cp865 | cp437 | cp437de | applemac | cp1252 | cp1250 | cp1257 | next⟩ 101 ⟨∗latin1 | decmulti | latin3 | latin5 | cp850 | cp858 | cp852 | cp1252 | next⟩

102 \ProvideTextCommandDefault{\textonehalf}{\ensuremath{\frac12}}

103 ⟨/latin1 | decmulti | latin3 | latin5 | cp850 | cp858 | cp852 | cp1252 | next⟩ 104 ⟨∗latin1 | decmulti | latin5 | cp850 | cp858 | cp852 | cp1252 | next⟩

105 \ProvideTextCommandDefault{\textonequarter}{\ensuremath{\frac14}}

106 ⟨/latin1 | decmulti | latin5 | cp850 | cp858 | cp852 | cp1252 | next⟩ 107 ⟨∗latin1 | latin5 | cp850 | cp858 | cp852 | cp1252 | next⟩

108 \ProvideTextCommandDefault{\textthreequarters}{\ensuremath{\frac34}}

109 ⟨/latin1 | latin5 | cp850 | cp858 | cp852 | cp1252 | next⟩ 110 ⟨∗applemac | cp850 | cp858 | cp865 | cp437 | cp437de | cp1252 | next⟩

111 \ProvideTextCommandDefault{\textflorin}{\textit{f}}

112 ⟨/applemac | cp850 | cp858 | cp865 | cp437 | cp437de | cp1252 | next⟩ 113 ⟨∗cp865 | cp437 | cp437de⟩

114 \ProvideTextCommandDefault{\textpeseta}{Pt}

115 ⟨/cp865 | cp437 | cp437de⟩ 116 ⟨∗cp850 | cp858 | cp852 | cp865 | cp437 | cp437de⟩

117 \ProvideTextCommandDefault{\textblacksquare}
118 {\vrule \@width .3em \@height .4em \@depth -.1em\relax}

119 ⟨/cp850 | cp858 | cp852 | cp865 | cp437 | cp437de⟩

Some commands can’t be faked, so we have them generate an error message.

120 ⟨∗latin1 | decmulti | latin5 | latin9 | cp850 | cp858 | cp865 | cp437 | cp437de⟩

121 \ProvideTextCommandDefault{\textcent}
122 {\TextSymbolUnavailable\textcent}
123 \ProvideTextCommandDefault{\textyen}
124 {\TextSymbolUnavailable\textyen}

125 ⟨/latin1 | decmulti | latin5 | latin9 | cp850 | cp858 | cp865 | cp437 | cp437de⟩ 126 ⟨∗applemac | cp1252 | next⟩

127 \ProvideTextCommandDefault{\textcent}
128 {\TextSymbolUnavailable\textcent}
129 \ProvideTextCommandDefault{\textyen}
130 {\TextSymbolUnavailable\textyen}

131 ⟨/applemac | cp1252 | next⟩ 132 ⟨∗latin9 | cp1252 | latin10⟩

133 \ProvideTextCommandDefault{\texteuro}
134 {\TextSymbolUnavailable\texteuro}

135 ⟨/latin9 | cp1252 | latin10⟩ 136 ⟨∗latin1 | decmulti | latin2 | latin3 | latin4 | latin5 | cp850 | cp858 | cp852 | cp865⟩

137 \ProvideTextCommandDefault{\textcurrency}
138 {\TextSymbolUnavailable\textcurrency}

139 ⟨/latin1 | decmulti | latin2 | latin3 | latin4 | latin5 | cp850 | cp858 | cp852 | cp865⟩ 140 ⟨∗applemac | cp1252 | cp1250 | cp1257 | next⟩

141 \ProvideTextCommandDefault{\textcurrency}
142 {\TextSymbolUnavailable\textcurrency}

143 ⟨/applemac | cp1252 | cp1250 | cp1257 | next⟩ 144 ⟨∗latin1 | latin5 | cp850 | cp858 | cp852 | cp1252 | cp1250⟩

145 \ProvideTextCommandDefault{\textbrokenbar}
146 {\TextSymbolUnavailable\textbrokenbar}

147 ⟨/latin1 | latin5 | cp850 | cp858 | cp852 | cp1252 | cp1250⟩ 148 ⟨∗latin3⟩

149 \ProvideTextCommandDefault{\textmalteseH}
150 {\TextSymbolUnavailable\textmalteseH}
151 \ProvideTextCommandDefault{\textmalteseh}
152 {\TextSymbolUnavailable\textmalteseh}

153 ⟨/latin3⟩ 154 ⟨∗latin4⟩

155 \ProvideTextCommandDefault{\textkra}
156 {\TextSymbolUnavailable\textkra}
157 \ProvideTextCommandDefault{\textTstroke}
158 {\TextSymbolUnavailable\textTstroke}
159 \ProvideTextCommandDefault{\texttstroke}
160 {\TextSymbolUnavailable\texttstroke}

161 ⟨/latin4⟩ 162 ⟨∗next⟩

163 \ProvideTextCommandDefault{\textquotesingle}
164 {\TextSymbolUnavailable\textquotesingle}
165 \ProvideTextCommandDefault{\capitalogonek}
166 {\TextSymbolUnavailable\capitalogonek}

167 ⟨/next⟩ 168 ⟨∗cp1250 | cp1252 | applemac | next⟩

169 \ProvideTextCommandDefault{\textperthousand}
170 {\TextSymbolUnavailable\textperthousand}

171 ⟨/cp1250 | cp1252 | applemac | next⟩ 172 ⟨∗applemacce⟩

173 \ProvideTextCommandDefault{\textdiv}
174 {\TextSymbolUnavailable\textdiv}

175 ⟨/applemacce⟩

Characters that are supposed to be used only in math will be defined by \providecommand because L^{A}TEX 2ε assumes that the font encoding for math fonts is static.

176 ⟨∗latin1 | decmulti | latin5 | latin9 | cp850 | cp858 | cp1252 | next⟩

177 \providecommand{\mathonesuperior}{{^1}}

178 ⟨/latin1 | decmulti | latin5 | latin9 | cp850 | cp858 | cp1252 | next⟩ 179 ⟨∗latin1 | decmulti | latin3 | latin5 | latin9 | cp850 | cp858 | cp1252 | next⟩

180 \providecommand{\maththreesuperior}{{^3}}

181 ⟨/latin1 | decmulti | latin3 | latin5 | latin9 | cp850 | cp858 | cp1252 | next⟩ 182 ⟨∗latin1 | decmulti | latin3 | latin5 | latin9 | cp850 | cp858 | cp865 | cp437 | cp437de⟩

183 \providecommand{\mathtwosuperior}{{^2}}

184 ⟨/latin1 | decmulti | latin3 | latin5 | latin9 | cp850 | cp858 | cp865 | cp437 | cp437de⟩ 185 ⟨∗cp1252 | next⟩

186 \providecommand{\mathtwosuperior}{{^2}}

187 ⟨/cp1252 | next⟩ 188 ⟨∗cp865 | cp437 | cp437de⟩

189 \providecommand{\mathnsuperior}{{^n}}

190 ⟨/cp865 | cp437 | cp437de⟩

5 The ASCII encoding

The ASCII encoding only allows characters in the range 32–127, so we only need to provide an empty .def file.

6 The ISO Latin-2 encoding

The ISO Latin-2 encoding file defines the characters in the ISO 8859-2 encoding. It was contributed by Petr Sojka (sojka@muni.cz) with small technical updates by Frank Mittelbach.

191 ⟨∗latin2⟩

192 \DeclareInputText{160}{\nobreakspace}
193 \DeclareInputText{176}{\textdegree}
194 \DeclareInputText{161}{\k A}
195 \DeclareInputText{177}{\k a}
196 \DeclareInputText{162}{\u{}}
197 \DeclareInputText{178}{\k\ }
198 \DeclareInputText{163}{\L}
199 \DeclareInputText{179}{\l}
200 \DeclareInputText{164}{\textcurrency}
201 \DeclareInputText{180}{\@tabacckludge'{}}
202 \DeclareInputText{165}{\v L}
203 \DeclareInputText{181}{\v l}
204 \DeclareInputText{166}{\@tabacckludge'S}
205 \DeclareInputText{182}{\@tabacckludge's}
206 \DeclareInputText{167}{\S}
207 \DeclareInputText{183}{\v{}}
208 \DeclareInputText{168}{\"{}}
209 \DeclareInputText{184}{\c\ }
210 \DeclareInputText{169}{\v S}
211 \DeclareInputText{185}{\v s}
212 \DeclareInputText{170}{\c S}
213 \DeclareInputText{186}{\c s}
214 \DeclareInputText{171}{\v T}
215 \DeclareInputText{187}{\v t}
216 \DeclareInputText{172}{\@tabacckludge'Z}
217 \DeclareInputText{188}{\@tabacckludge'z}
218 \DeclareInputText{173}{\-}
219 \DeclareInputText{189}{\H{}}
220 \DeclareInputText{174}{\v Z}
221 \DeclareInputText{190}{\v z}
222 \DeclareInputText{175}{\.Z}
223 \DeclareInputText{191}{\.z}

224 \DeclareInputText{192}{\@tabacckludge'R}
225 \DeclareInputText{208}{\DJ}
226 \DeclareInputText{193}{\@tabacckludge'A}
227 \DeclareInputText{209}{\@tabacckludge'N}
228 \DeclareInputText{194}{\^A}
229 \DeclareInputText{210}{\v N}
230 \DeclareInputText{195}{\u A}
231 \DeclareInputText{211}{\@tabacckludge'O}
232 \DeclareInputText{196}{\"A}
233 \DeclareInputText{212}{\^O}
234 \DeclareInputText{197}{\@tabacckludge'L}
235 \DeclareInputText{213}{\H O}
236 \DeclareInputText{198}{\@tabacckludge'C}
237 \DeclareInputText{214}{\"O}
238 \DeclareInputText{199}{\c C}
239 \DeclareInputMath{215}{\times}
240 \DeclareInputText{200}{\v C}
241 \DeclareInputText{216}{\v R}
242 \DeclareInputText{201}{\@tabacckludge'E}
243 \DeclareInputText{217}{\r U}
244 \DeclareInputText{202}{\k E}
245 \DeclareInputText{218}{\@tabacckludge'U}
246 \DeclareInputText{203}{\"E}
247 \DeclareInputText{219}{\H U}
248 \DeclareInputText{204}{\v E}
249 \DeclareInputText{220}{\"U}
250 \DeclareInputText{205}{\@tabacckludge'I}
251 \DeclareInputText{221}{\@tabacckludge'Y}
252 \DeclareInputText{206}{\^I}
253 \DeclareInputText{222}{\c T}
254 \DeclareInputText{207}{\v D}
255 \DeclareInputText{223}{\ss}
256 \DeclareInputText{224}{\@tabacckludge'r}
257 \DeclareInputText{240}{\dj}
258 \DeclareInputText{225}{\@tabacckludge'a}
259 \DeclareInputText{241}{\@tabacckludge'n}
260 \DeclareInputText{226}{\^a}
261 \DeclareInputText{242}{\v n}
262 \DeclareInputText{227}{\u a}
263 \DeclareInputText{243}{\@tabacckludge'o}
264 \DeclareInputText{228}{\"a}
265 \DeclareInputText{244}{\^o}
266 \DeclareInputText{229}{\@tabacckludge'l}
267 \DeclareInputText{245}{\H o}
268 \DeclareInputText{230}{\@tabacckludge'c}
269 \DeclareInputText{246}{\"o}
270 \DeclareInputText{231}{\c c}
271 \DeclareInputMath{247}{\div}
272 \DeclareInputText{232}{\v c}
273 \DeclareInputText{248}{\v r}
274 \DeclareInputText{233}{\@tabacckludge'e}
275 \DeclareInputText{249}{\r u}
276 \DeclareInputText{234}{\k e}
277 \DeclareInputText{250}{\@tabacckludge'u}

278 \DeclareInputText{235}{\"e}
279 \DeclareInputText{251}{\H u}
280 \DeclareInputText{236}{\v e}
281 \DeclareInputText{252}{\"u}
282 \DeclareInputText{237}{\@tabacckludge'\i}
283 \DeclareInputText{253}{\@tabacckludge'y}
284 \DeclareInputText{238}{\^\i}
285 \DeclareInputText{254}{\c t}
286 \DeclareInputText{239}{\v d}
287 \DeclareInputText{255}{\.{}}

288 ⟨/latin2⟩

7 The ISO Latin-3 encoding

The ISO Latin-3 encoding file defines the characters in the ISO 8859-3 encoding. It was contributed by by Jörg Knappen (joerg.knappen@uni-mainz.de) and adapted by Chris Rowley.

It can be used for general purpose applications in typical office environments in the following languages: Afrikaans, Catalan, English, Esperanto, French, Gallician, German, Italian, Maltese, and Turkish.

289 ⟨∗latin3⟩

290 \DeclareInputText{160}{\nobreakspace}
291 \DeclareInputText{176}{\textdegree}
292 \DeclareInputText{161}{\textmalteseH}
293 \DeclareInputText{177}{\textmalteseh}
294 \DeclareInputText{162}{\u{}}
295 \DeclareInputMath{178}{\mathtwosuperior}
296 \DeclareInputText{163}{\pounds}
297 \DeclareInputMath{179}{\maththreesuperior}
298 \DeclareInputText{164}{\textcurrency}
299 \DeclareInputText{180}{\@tabacckludge'{}}
300 % \DeclareInputText{165}{\notdef}
301 \DeclareInputMath{181}{\mu}
302 \DeclareInputText{166}{\^H}
303 % NOT: \DeclareInputText{182}{h\llap{\^{}}} % \^h would be too tall
304 \DeclareInputText{182}{\^h}
305 \DeclareInputText{167}{\S}
306 \DeclareInputText{183}{\textperiodcentered}
307 \DeclareInputText{168}{\"{}}
308 \DeclareInputText{184}{\c\ }
309 \DeclareInputText{169}{\.I}
310 \DeclareInputText{185}{\i}
311 \DeclareInputText{170}{\c S}
312 \DeclareInputText{186}{\c s}
313 \DeclareInputText{171}{\u G}
314 \DeclareInputText{187}{\u g}
315 \DeclareInputText{172}{\^J}
316 \DeclareInputText{188}{\^\j}
317 \DeclareInputText{173}{\-}
318 \DeclareInputText{189}{\textonehalf}
319 % \DeclareInputText{174}{\notdef}
320 % \DeclareInputText{190}{\notdef}

321 \DeclareInputText{175}{\.Z}
322 \DeclareInputText{191}{\.z}
323 \DeclareInputText{192}{\@tabacckludge`A}
324 % \DeclareInputText{208}{\notdef}
325 \DeclareInputText{193}{\@tabacckludge'A}
326 \DeclareInputText{209}{\~N}
327 \DeclareInputText{194}{\^A}
328 \DeclareInputText{210}{\@tabacckludge`O}
329 % \DeclareInputText{195}{\notdef}
330 \DeclareInputText{211}{\@tabacckludge'O}
331 \DeclareInputText{196}{\"A}
332 \DeclareInputText{212}{\^O}
333 \DeclareInputText{197}{\.C}
334 \DeclareInputText{213}{\.G}
335 \DeclareInputText{198}{\^C}
336 \DeclareInputText{214}{\"O}
337 \DeclareInputText{199}{\c C}
338 \DeclareInputMath{215}{\times}
339 \DeclareInputText{200}{\@tabacckludge`E}
340 \DeclareInputText{216}{\^G}
341 \DeclareInputText{201}{\@tabacckludge'E}
342 \DeclareInputText{217}{\@tabacckludge`U}
343 \DeclareInputText{202}{\^E}
344 \DeclareInputText{218}{\@tabacckludge'U}
345 \DeclareInputText{203}{\"E}
346 \DeclareInputText{219}{\^U}
347 \DeclareInputText{204}{\@tabacckludge`I}
348 \DeclareInputText{220}{\"U}
349 \DeclareInputText{205}{\@tabacckludge'I}
350 \DeclareInputText{221}{\u U}
351 \DeclareInputText{206}{\^I}
352 \DeclareInputText{222}{\^S}
353 \DeclareInputText{207}{\"I}
354 \DeclareInputText{223}{\ss}
355 \DeclareInputText{224}{\@tabacckludge`a}
356 % \DeclareInputText{240}{\notdef}
357 \DeclareInputText{225}{\@tabacckludge'a}
358 \DeclareInputText{241}{\~n}
359 \DeclareInputText{226}{\^a}
360 \DeclareInputText{242}{\@tabacckludge`o}
361 % \DeclareInputText{227}{\notdef}
362 \DeclareInputText{243}{\@tabacckludge'o}
363 \DeclareInputText{228}{\"a}
364 \DeclareInputText{244}{\^o}
365 \DeclareInputText{229}{\.c}
366 \DeclareInputText{245}{\.g}
367 \DeclareInputText{230}{\^c}
368 \DeclareInputText{246}{\"o}
369 \DeclareInputText{231}{\c c}
370 \DeclareInputMath{247}{\div}
371 \DeclareInputText{232}{\@tabacckludge`e}
372 \DeclareInputText{248}{\^g}
373 \DeclareInputText{233}{\@tabacckludge'e}

374 \DeclareInputText{249}{\@tabacckludge`u}
375 \DeclareInputText{234}{\^e}
376 \DeclareInputText{250}{\@tabacckludge'u}
377 \DeclareInputText{235}{\"e}
378 \DeclareInputText{251}{\^u}
379 \DeclareInputText{236}{\@tabacckludge`\i}
380 \DeclareInputText{252}{\"u}
381 \DeclareInputText{237}{\@tabacckludge'\i}
382 \DeclareInputText{253}{\u u}
383 \DeclareInputText{238}{\^\i}
384 \DeclareInputText{254}{\^s}
385 \DeclareInputText{239}{\"\i}
386 \DeclareInputText{255}{\.{}}

387 ⟨/latin3⟩

8 The ISO Latin-4 encoding

The ISO Latin-4 encoding file defines the characters in the ISO 8859-4 encoding. It was contributed by Hana Skoumalová (hana.skoumalova@ff.cuni.cz).

It was created for Estonian, Latvian, Lithuanian, Finnish, Lappish, Swedish, Norvegian, Danish, Icelandic and Greenlandic Inuit. Some alphabets, however, are incomplete (Greenlandic, Icelandic and Lappish). Some glyphs are not available in the fonts. For example, the Greenlandic character ‘kra’ is not available at all; the Latvian and Lithuanian characters not contained in other encodings are constructed from components and they do not look good.

388 ⟨∗latin4⟩

389 \DeclareInputText{160}{\nobreakspace}
390 \DeclareInputText{161}{\k A}
391 \DeclareInputText{162}{\textkra}%% Greenlandic Inuit
392 \DeclareInputText{163}{\c R}
393 \DeclareInputText{164}{\textcurrency}
394 \DeclareInputText{165}{\~I}
395 \DeclareInputText{166}{\c L}
396 \DeclareInputText{167}{\S}
397 \DeclareInputText{168}{\"\ }
398 \DeclareInputText{169}{\v S}
399 \DeclareInputText{170}{\@tabacckludge=E}
400 \DeclareInputText{171}{\c G}
401 \DeclareInputText{172}{\textTstroke}%% Northern Sami
402 \DeclareInputText{173}{\-}
403 \DeclareInputText{174}{\v Z}
404 \DeclareInputText{175}{\@tabacckludge={}}
405 \DeclareInputText{176}{\textdegree}
406 \DeclareInputText{177}{\k a}
407 \DeclareInputText{178}{\k\ }
408 \DeclareInputText{179}{\c r}
409 \DeclareInputText{180}{\@tabacckludge'{}}
410 \DeclareInputText{181}{\~\i}
411 \DeclareInputText{182}{\c l}
412 \DeclareInputText{183}{\v\ }
413 \DeclareInputText{184}{\c\ }
414 \DeclareInputText{185}{\v s}

415 \DeclareInputText{186}{\@tabacckludge=e}
416 \DeclareInputText{187}{\c g}
417 \DeclareInputText{188}{\texttstroke}%% Northern Sami
418 \DeclareInputText{189}{\NG}
419 \DeclareInputText{190}{\v z}
420 \DeclareInputText{191}{\ng}
421 \DeclareInputText{192}{\@tabacckludge=A}
422 \DeclareInputText{193}{\@tabacckludge'A}
423 \DeclareInputText{194}{\^A}
424 \DeclareInputText{195}{\~A}
425 \DeclareInputText{196}{\"A}
426 \DeclareInputText{197}{\r A}
427 \DeclareInputText{198}{\AE}
428 \DeclareInputText{199}{\k I}
429 \DeclareInputText{200}{\v C}
430 \DeclareInputText{201}{\@tabacckludge'E}
431 \DeclareInputText{202}{\k E}
432 \DeclareInputText{203}{\"E}
433 \DeclareInputText{204}{\.{E}}
434 \DeclareInputText{205}{\@tabacckludge'I}
435 \DeclareInputText{206}{\^I}
436 \DeclareInputText{207}{\@tabacckludge=I}
437 \DeclareInputText{208}{\DJ}
438 \DeclareInputText{209}{\c N}
439 \DeclareInputText{210}{\@tabacckludge=O}
440 \DeclareInputText{211}{\c K}
441 \DeclareInputText{212}{\^O}
442 \DeclareInputText{213}{\~O}
443 \DeclareInputText{214}{\"O}
444 \DeclareInputMath{215}{\times}
445 \DeclareInputText{216}{\O}
446 \DeclareInputText{217}{\k U}
447 \DeclareInputText{218}{\@tabacckludge'U}
448 \DeclareInputText{219}{\^U}
449 \DeclareInputText{220}{\"U}
450 \DeclareInputText{221}{\~U}
451 \DeclareInputText{222}{\@tabacckludge=U}
452 \DeclareInputText{223}{\ss}
453 \DeclareInputText{224}{\@tabacckludge=a}
454 \DeclareInputText{225}{\@tabacckludge'a}
455 \DeclareInputText{226}{\^a}
456 \DeclareInputText{227}{\~a}
457 \DeclareInputText{228}{\"a}
458 \DeclareInputText{229}{\r a}
459 \DeclareInputText{230}{\ae}
460 \DeclareInputText{231}{\k i}
461 \DeclareInputText{232}{\v c}
462 \DeclareInputText{233}{\@tabacckludge'e}
463 \DeclareInputText{234}{\k e}
464 \DeclareInputText{235}{\"e}
465 \DeclareInputText{236}{\.{e}}
466 \DeclareInputText{237}{\@tabacckludge'\i}
467 \DeclareInputText{238}{\^\i}

468 \DeclareInputText{239}{\@tabacckludge=\i}
469 \DeclareInputText{240}{\dj}
470 \DeclareInputText{241}{\c n}
471 \DeclareInputText{242}{\@tabacckludge=o}
472 \DeclareInputText{243}{\c k}
473 \DeclareInputText{244}{\^o}
474 \DeclareInputText{245}{\~o}
475 \DeclareInputText{246}{\"o}
476 \DeclareInputMath{247}{\div}
477 \DeclareInputText{248}{\o}
478 \DeclareInputText{249}{\k u}
479 \DeclareInputText{250}{\@tabacckludge'u}
480 \DeclareInputText{251}{\^u}
481 \DeclareInputText{252}{\"u}
482 \DeclareInputText{253}{\~u}
483 \DeclareInputText{254}{\@tabacckludge=u}
484 \DeclareInputText{255}{\.\ }

485 ⟨/latin4⟩

9 The ISO Latin-5 encoding

The ISO Latin-5 encoding file defines the characters in the ISO 8859-9 encoding, which describes Latin Alphabet No. 5. It was contributed by H. Turgut Uyar (uyar@cs.itu.edu.tr); it is used for Turkish.

486 ⟨∗latin5⟩

487 \DeclareInputText{160}{\nobreakspace}
488 \DeclareInputText{176}{\textdegree}
489 \DeclareInputText{161}{\textexclamdown}
490 \DeclareInputMath{177}{\pm}
491 \DeclareInputText{162}{\textcent}
492 \DeclareInputMath{178}{\mathtwosuperior}
493 \DeclareInputText{163}{\pounds}
494 \DeclareInputMath{179}{\maththreesuperior}
495 \DeclareInputText{164}{\textcurrency}
496 \DeclareInputText{180}{\@tabacckludge'{}}
497 \DeclareInputText{165}{\textyen}
498 \DeclareInputMath{181}{\mu}
499 \DeclareInputText{166}{\textbrokenbar}
500 \DeclareInputText{182}{\P}
501 \DeclareInputText{167}{\S}
502 \DeclareInputText{183}{\textperiodcentered}
503 \DeclareInputText{168}{\"{}}
504 \DeclareInputText{184}{\c\ }
505 \DeclareInputText{169}{\copyright}
506 \DeclareInputMath{185}{\mathonesuperior}
507 \DeclareInputText{170}{\textordfeminine}
508 \DeclareInputText{186}{\textordmasculine}
509 \DeclareInputText{171}{\guillemotleft}
510 \DeclareInputText{187}{\guillemotright}
511 \DeclareInputMath{172}{\lnot}
512 \DeclareInputText{188}{\textonequarter}
513 \DeclareInputText{173}{\-}
514 \DeclareInputText{189}{\textonehalf}

515 \DeclareInputText{174}{\textregistered}
516 \DeclareInputText{190}{\textthreequarters}
517 \DeclareInputText{175}{\@tabacckludge={}}
518 \DeclareInputText{191}{\textquestiondown}
519 \DeclareInputText{192}{\@tabacckludge`A}
520 \DeclareInputText{208}{\u G}
521 \DeclareInputText{193}{\@tabacckludge'A}
522 \DeclareInputText{209}{\~N}
523 \DeclareInputText{194}{\^A}
524 \DeclareInputText{210}{\@tabacckludge`O}
525 \DeclareInputText{195}{\~A}
526 \DeclareInputText{211}{\@tabacckludge'O}
527 \DeclareInputText{196}{\"A}
528 \DeclareInputText{212}{\^O}
529 \DeclareInputText{197}{\r A}
530 \DeclareInputText{213}{\~O}
531 \DeclareInputText{198}{\AE}
532 \DeclareInputText{214}{\"O}
533 \DeclareInputText{199}{\c C}
534 \DeclareInputMath{215}{\times}
535 \DeclareInputText{200}{\@tabacckludge`E}
536 \DeclareInputText{216}{\O}
537 \DeclareInputText{201}{\@tabacckludge'E}
538 \DeclareInputText{217}{\@tabacckludge`U}
539 \DeclareInputText{202}{\^E}
540 \DeclareInputText{218}{\@tabacckludge'U}
541 \DeclareInputText{203}{\"E}
542 \DeclareInputText{219}{\^U}
543 \DeclareInputText{204}{\@tabacckludge`I}
544 \DeclareInputText{220}{\"U}
545 \DeclareInputText{205}{\@tabacckludge'I}
546 \DeclareInputText{221}{\.I}
547 \DeclareInputText{206}{\^I}
548 \DeclareInputText{222}{\c S}
549 \DeclareInputText{207}{\"I}
550 \DeclareInputText{223}{\ss}
551 \DeclareInputText{224}{\@tabacckludge`a}
552 \DeclareInputText{240}{\u g}
553 \DeclareInputText{225}{\@tabacckludge'a}
554 \DeclareInputText{241}{\~n}
555 \DeclareInputText{226}{\^a}
556 \DeclareInputText{242}{\@tabacckludge`o}
557 \DeclareInputText{227}{\~a}
558 \DeclareInputText{243}{\@tabacckludge'o}
559 \DeclareInputText{228}{\"a}
560 \DeclareInputText{244}{\^o}
561 \DeclareInputText{229}{\r a}
562 \DeclareInputText{245}{\~o}
563 \DeclareInputText{230}{\ae}
564 \DeclareInputText{246}{\"o}
565 \DeclareInputText{231}{\c c}
566 \DeclareInputMath{247}{\div}
567 \DeclareInputText{232}{\@tabacckludge`e}

568 \DeclareInputText{248}{\o}
569 \DeclareInputText{233}{\@tabacckludge'e}
570 \DeclareInputText{249}{\@tabacckludge`u}
571 \DeclareInputText{234}{\^e}
572 \DeclareInputText{250}{\@tabacckludge'u}
573 \DeclareInputText{235}{\"e}
574 \DeclareInputText{251}{\^u}
575 \DeclareInputText{236}{\@tabacckludge`\i}
576 \DeclareInputText{252}{\"u}
577 \DeclareInputText{237}{\@tabacckludge'\i}
578 \DeclareInputText{253}{\i }
579 \DeclareInputText{238}{\^\i}
580 \DeclareInputText{254}{\c s}
581 \DeclareInputText{239}{\"\i}
582 \DeclareInputText{255}{\"y}

583 ⟨/latin5⟩

10 DEC Multinational Character Set

The DECMultinational character set, used by the OpenVMS operating system, is slightly different from the ISO Latin 1 character set.

Reference: Digital Equipment Corporation VT330/VT340 Programmer Reference Manual, Volume 1: Text Programming, page 22.

This encoding was provided by M.Y. Chartoire IPNL-IN2P3

 m.chartoire@ipnl.in2p3.fr

584 ⟨∗decmulti⟩

585 \DeclareInputText{176}{\textdegree}
586 \DeclareInputText{161}{\textexclamdown}
587 \DeclareInputMath{177}{\pm}
588 \DeclareInputText{162}{\textcent}
589 \DeclareInputMath{178}{\mathtwosuperior}
590 \DeclareInputText{163}{\pounds}
591 \DeclareInputMath{179}{\maththreesuperior}
592 \DeclareInputText{165}{\textyen}
593 \DeclareInputMath{181}{\mu}
594 \DeclareInputText{182}{\P}
595 \DeclareInputText{167}{\S}
596 \DeclareInputText{183}{\textperiodcentered}
597 \DeclareInputText{168}{\textcurrency}
598 \DeclareInputText{169}{\copyright}
599 \DeclareInputMath{185}{\mathonesuperior}
600 \DeclareInputText{170}{\textordfeminine}
601 \DeclareInputText{186}{\textordmasculine}
602 \DeclareInputText{171}{\guillemotleft}
603 \DeclareInputText{187}{\guillemotright}
604 \DeclareInputText{188}{\textonequarter}
605 \DeclareInputText{189}{\textonehalf}
606 \DeclareInputText{191}{\textquestiondown}
607 \DeclareInputText{192}{\@tabacckludge`A}
608 \DeclareInputText{193}{\@tabacckludge'A}
609 \DeclareInputText{209}{\~N}
610 \DeclareInputText{194}{\^A}

611 \DeclareInputText{210}{\@tabacckludge`O}
612 \DeclareInputText{195}{\~A}
613 \DeclareInputText{211}{\@tabacckludge'O}
614 \DeclareInputText{196}{\"A}
615 \DeclareInputText{212}{\^O}
616 \DeclareInputText{197}{\r A}
617 \DeclareInputText{213}{\~O}
618 \DeclareInputText{198}{\AE}
619 \DeclareInputText{214}{\"O}
620 \DeclareInputText{199}{\c C}
621 \DeclareInputText{215}{\OE}
622 \DeclareInputText{200}{\@tabacckludge`E}
623 \DeclareInputText{216}{\O}
624 \DeclareInputText{201}{\@tabacckludge'E}
625 \DeclareInputText{217}{\@tabacckludge`U}
626 \DeclareInputText{202}{\^E}
627 \DeclareInputText{218}{\@tabacckludge'U}
628 \DeclareInputText{203}{\"E}
629 \DeclareInputText{219}{\^U}
630 \DeclareInputText{204}{\@tabacckludge`I}
631 \DeclareInputText{220}{\"U}
632 \DeclareInputText{205}{\@tabacckludge'I}
633 \DeclareInputText{221}{\"Y}
634 \DeclareInputText{206}{\^I}
635 \DeclareInputText{207}{\"I}
636 \DeclareInputText{223}{\ss}
637 \DeclareInputText{224}{\@tabacckludge`a}
638 \DeclareInputText{225}{\@tabacckludge'a}
639 \DeclareInputText{241}{\~n}
640 \DeclareInputText{226}{\^a}
641 \DeclareInputText{242}{\@tabacckludge`o}
642 \DeclareInputText{227}{\~a}
643 \DeclareInputText{243}{\@tabacckludge'o}
644 \DeclareInputText{228}{\"a}
645 \DeclareInputText{244}{\^o}
646 \DeclareInputText{229}{\r a}
647 \DeclareInputText{245}{\~o}
648 \DeclareInputText{230}{\ae}
649 \DeclareInputText{246}{\"o}
650 \DeclareInputText{231}{\c c}
651 \DeclareInputText{247}{\oe}
652 \DeclareInputText{232}{\@tabacckludge`e}
653 \DeclareInputText{248}{\o}
654 \DeclareInputText{233}{\@tabacckludge'e}
655 \DeclareInputText{249}{\@tabacckludge`u}
656 \DeclareInputText{234}{\^e}
657 \DeclareInputText{250}{\@tabacckludge'u}
658 \DeclareInputText{235}{\"e}
659 \DeclareInputText{251}{\^u}
660 \DeclareInputText{236}{\@tabacckludge`\i}
661 \DeclareInputText{252}{\"u}
662 \DeclareInputText{237}{\@tabacckludge'\i}
663 \DeclareInputText{253}{\"y}
664 \DeclareInputText{238}{\^\i}

665 \DeclareInputText{239}{\"\i}

666 ⟨/decmulti⟩

11 The IBM code pages 850 and 858

This input encoding was contributed by Timo Knuutila (knuutila@cs.utu.fi), and edited by Christian Bartels (ii140ba@vm1.rz.rwth-aachen.de).

The DOS graphics ‘letters’ and a few other positions are ignored (left undefined).

The 858 code page is identical to the 850 except that

667 ⟨∗cp850 | cp858⟩

668 \DeclareInputText{128}{\c C}
669 \DeclareInputText{144}{\@tabacckludge'E}
670 \DeclareInputText{129}{\"u}
671 \DeclareInputText{145}{\ae}
672 \DeclareInputText{130}{\@tabacckludge'e}
673 \DeclareInputText{146}{\AE}
674 \DeclareInputText{131}{\^a}
675 \DeclareInputText{147}{\^o}
676 \DeclareInputText{132}{\"a}
677 \DeclareInputText{148}{\"o}
678 \DeclareInputText{133}{\@tabacckludge`a}
679 \DeclareInputText{149}{\@tabacckludge`o}
680 \DeclareInputText{134}{\r a}
681 \DeclareInputText{150}{\^u}
682 \DeclareInputText{135}{\c c}
683 \DeclareInputText{151}{\@tabacckludge`u}
684 \DeclareInputText{136}{\^e}
685 \DeclareInputText{152}{\"y}
686 \DeclareInputText{137}{\"e}
687 \DeclareInputText{153}{\"O}
688 \DeclareInputText{138}{\@tabacckludge`e}
689 \DeclareInputText{154}{\"U}
690 \DeclareInputText{139}{\"\i}
691 \DeclareInputText{155}{\o}
692 \DeclareInputText{140}{\^\i}
693 \DeclareInputText{156}{\pounds}
694 \DeclareInputText{141}{\@tabacckludge`\i}
695 \DeclareInputText{157}{\O}
696 \DeclareInputText{142}{\"A}
697 \DeclareInputMath{158}{\times}
698 \DeclareInputText{143}{\r A}
699 \DeclareInputText{159}{\textflorin}
700 \DeclareInputText{160}{\@tabacckludge'a}
701 \DeclareInputText{161}{\@tabacckludge'\i}
702 \DeclareInputText{162}{\@tabacckludge'o}
703 \DeclareInputText{163}{\@tabacckludge'u}
704 \DeclareInputText{164}{\~n}               %% "B0-"B4: DG
705 \DeclareInputText{165}{\~N}
706 \DeclareInputText{181}{\@tabacckludge'A}
707 \DeclareInputText{166}{\textordfeminine}
708 \DeclareInputText{182}{\^A}

709 \DeclareInputText{167}{\textordmasculine}
710 \DeclareInputText{183}{\@tabacckludge`A}
711 \DeclareInputText{168}{\textquestiondown}
712 \DeclareInputText{184}{\copyright}
713 \DeclareInputText{169}{\textregistered}
714 \DeclareInputMath{170}{\lnot}
715 \DeclareInputText{171}{\textonehalf}
716 \DeclareInputText{172}{\textonequarter}   %% "B9-"BC: DG
717 \DeclareInputText{173}{\textexclamdown}
718 \DeclareInputText{189}{\textcent}
719 \DeclareInputText{174}{\guillemotleft}
720 \DeclareInputText{190}{\textyen}
721 \DeclareInputText{175}{\guillemotright}
722 % "BF-"C5: DG
723 \DeclareInputText{208}{\dh}
724 \DeclareInputText{209}{\DH}
725 \DeclareInputText{210}{\^E}
726 \DeclareInputText{211}{\"E}
727 \DeclareInputText{212}{\@tabacckludge`E}

Here is the only point in which the two code pages differ!

728 ⟨−cp858⟩\DeclareInputText{213}{\i}
729 ⟨−cp850⟩\DeclareInputText{213}{\texteuro}
730 \DeclareInputText{198}{\~a}
731 \DeclareInputText{214}{\@tabacckludge'I}
732 \DeclareInputText{199}{\~A}
733 \DeclareInputText{215}{\^I}
734 \DeclareInputText{216}{\"I}
735 % "C8-"CE: DG
736 % "D9-"DC: DG
737 \DeclareInputText{221}{\textbrokenbar}
738 \DeclareInputText{222}{\@tabacckludge`I}
739 \DeclareInputText{207}{\textcurrency}  %% "DF: DG
740 \DeclareInputText{224}{\@tabacckludge'O}
741 \DeclareInputText{240}{\textendash}
742 \DeclareInputText{225}{\ss}
743 \DeclareInputMath{241}{\pm}
744 \DeclareInputText{226}{\^O}
745 % "F2: DG (not double underline, or equals?)
746 \DeclareInputText{227}{\@tabacckludge`O}
747 \DeclareInputText{243}{\textthreequarters}
748 \DeclareInputText{228}{\~o}
749 \DeclareInputText{244}{\P}
750 \DeclareInputText{229}{\~O}
751 \DeclareInputText{245}{\S}
752 \DeclareInputMath{230}{\mu}
753 \DeclareInputMath{246}{\div}
754 \DeclareInputText{231}{\th}
755 \DeclareInputText{247}{\c\ }
756 \DeclareInputText{232}{\TH}
757 \DeclareInputText{248}{\textdegree}
758 \DeclareInputText{233}{\@tabacckludge'U}
759 \DeclareInputText{249}{\"{}}

760 \DeclareInputText{234}{\^U}
761 \DeclareInputText{250}{\textperiodcentered}
762 \DeclareInputText{235}{\@tabacckludge`U}
763 \DeclareInputMath{251}{\mathonesuperior}
764 \DeclareInputText{236}{\@tabacckludge'y}
765 \DeclareInputMath{252}{\maththreesuperior}
766 \DeclareInputText{237}{\@tabacckludge'Y}
767 \DeclareInputMath{253}{\mathtwosuperior}
768 \DeclareInputText{238}{\@tabacckludge={}}
769 \DeclareInputText{254}{\textblacksquare} % right name?
770 \DeclareInputText{239}{\@tabacckludge'{}}
771 \DeclareInputText{255}{\nobreakspace}

772 ⟨/cp850 | cp858⟩

12 The IBM code page 852

This input encoding was contributed by Petr Sojka (sojka@Muni.cz).

773 ⟨∗cp852⟩

774 \DeclareInputText{128}{\c C}
775 \DeclareInputText{144}{\@tabacckludge'E}
776 \DeclareInputText{129}{\"u}
777 \DeclareInputText{145}{\@tabacckludge'L}
778 \DeclareInputText{130}{\@tabacckludge'e}
779 \DeclareInputText{146}{\@tabacckludge'l}
780 \DeclareInputText{131}{\^a}
781 \DeclareInputText{147}{\^o}
782 \DeclareInputText{132}{\"a}
783 \DeclareInputText{148}{\"o}
784 \DeclareInputText{133}{\r u}
785 \DeclareInputText{149}{\v L}
786 \DeclareInputText{134}{\@tabacckludge'c}
787 \DeclareInputText{150}{\v l}
788 \DeclareInputText{135}{\c c}
789 \DeclareInputText{151}{\@tabacckludge'S}
790 \DeclareInputText{136}{\l}
791 \DeclareInputText{152}{\@tabacckludge's}
792 \DeclareInputText{137}{\"e}
793 \DeclareInputText{153}{\"O}
794 \DeclareInputText{138}{\H O}
795 \DeclareInputText{154}{\"U}
796 \DeclareInputText{139}{\H o}
797 \DeclareInputText{155}{\v T}
798 \DeclareInputText{140}{\^\i}
799 \DeclareInputText{156}{\v t}
800 \DeclareInputText{141}{\@tabacckludge'Z}
801 \DeclareInputText{157}{\L}
802 \DeclareInputText{142}{\"A}
803 \DeclareInputMath{158}{\times}
804 \DeclareInputText{143}{\@tabacckludge'C}
805 \DeclareInputText{159}{\v c}
806 \DeclareInputText{160}{\@tabacckludge'a}
807 \DeclareInputText{161}{\@tabacckludge'\i}

808 \DeclareInputText{162}{\@tabacckludge'o}
809 \DeclareInputText{163}{\@tabacckludge'u}
810 \DeclareInputText{164}{\k A}
811 \DeclareInputText{165}{\k a}
812 \DeclareInputText{166}{\v Z}
813 \DeclareInputText{167}{\v z}
814 \DeclareInputText{168}{\k E}
815 \DeclareInputText{169}{\k e}
816 \DeclareInputText{170}{\nobreakspace}  % or empty??
817 \DeclareInputText{171}{\@tabacckludge'z}
818 \DeclareInputText{172}{\v C}
819 \DeclareInputText{173}{\c s}
820 \DeclareInputText{174}{\guillemotleft}
821 \DeclareInputText{175}{\guillemotright}
822 \DeclareInputText{181}{\@tabacckludge'A}
823 \DeclareInputText{182}{\^A}
824 \DeclareInputText{183}{\v E}
825 \DeclareInputText{184}{\c S}
826 \DeclareInputText{189}{\.Z}
827 \DeclareInputText{190}{\.z}
828 \DeclareInputText{198}{\u A}
829 \DeclareInputText{199}{\u a}
830 \DeclareInputText{207}{\textcurrency}
831 \DeclareInputText{208}{\dj}
832 \DeclareInputText{209}{\DJ}
833 \DeclareInputText{210}{\v D}
834 \DeclareInputText{211}{\"E}
835 \DeclareInputText{212}{\v d} % d caron
836 \DeclareInputText{213}{\v N}
837 \DeclareInputText{214}{\@tabacckludge'I}
838 \DeclareInputText{215}{\^I}
839 \DeclareInputText{216}{\v e}
840 \DeclareInputText{221}{\c T}
841 \DeclareInputText{222}{\r U}
842 \DeclareInputText{224}{\@tabacckludge'O}
843 \DeclareInputText{240}{\textendash}
844 \DeclareInputText{225}{\ss}
845 \DeclareInputText{241}{\H{}}
846 \DeclareInputText{226}{\^O}
847 \DeclareInputText{242}{\k\ }
848 \DeclareInputText{227}{\@tabacckludge'N}
849 \DeclareInputText{243}{\v{}}
850 \DeclareInputText{228}{\@tabacckludge'n}
851 \DeclareInputText{244}{\u{}}
852 \DeclareInputText{229}{\v n}
853 \DeclareInputText{245}{\S}
854 \DeclareInputText{230}{\v S}
855 \DeclareInputMath{246}{\div}
856 \DeclareInputText{231}{\v s}
857 \DeclareInputText{247}{\c\ }
858 \DeclareInputText{232}{\@tabacckludge'R}
859 \DeclareInputText{248}{\textdegree}
860 \DeclareInputText{233}{\@tabacckludge'U}

861 \DeclareInputText{249}{\"{}}
862 \DeclareInputText{234}{\@tabacckludge'r}
863 \DeclareInputText{250}{\.{}}
864 \DeclareInputText{235}{\H U}
865 \DeclareInputText{251}{\H u}
866 \DeclareInputText{236}{\@tabacckludge'y}
867 \DeclareInputText{252}{\v R}
868 \DeclareInputText{237}{\@tabacckludge'Y}
869 \DeclareInputText{253}{\v r}
870 \DeclareInputText{238}{\c t}
871 \DeclareInputText{254}{\textblacksquare} % right name?
872 \DeclareInputText{239}{\@tabacckludge'{}}
873 \DeclareInputText{255}{\nobreakspace}

874 ⟨/cp852⟩

13 The IBM code pages 437 and 865

This input encoding is based on work by

 Volker Kunert (volker@numsun1.mathematik.uni-halle.de)
 and bontus@al6000.physik.uni-siegen.de.

The changes for cp865 are based on work by Søren Sandmann (sandmann@ daimi.aau.dk), with thanks to them all.

The DOS graphics ‘letters’ and a few other positions are ignored (left undefined).

Unfortunately, in cp437 there is no agreement as to whether slot E1 should be ‘β’ or ‘ß’, so we provide two variants, one (cp437) with ‘β’ and one (cp437de) with ‘ß’.

875 ⟨∗cp437 | cp437de | cp865⟩

876 \DeclareInputText{128}{{\c C}}
877 \DeclareInputText{129}{\"u}
878 \DeclareInputText{130}{\@tabacckludge'e}
879 \DeclareInputText{131}{\^a}
880 \DeclareInputText{132}{\"a}
881 \DeclareInputText{133}{\@tabacckludge`a}
882 \DeclareInputText{134}{\r a}
883 \DeclareInputText{135}{{\c c}}
884 \DeclareInputText{136}{\^e}
885 \DeclareInputText{137}{\"e}
886 \DeclareInputText{138}{\@tabacckludge`e}
887 \DeclareInputText{139}{\"\i}
888 \DeclareInputText{140}{\^\i}
889 \DeclareInputText{141}{\@tabacckludge`\i}
890 \DeclareInputText{142}{\"A}
891 \DeclareInputText{143}{\r A}
892 \DeclareInputText{144}{\@tabacckludge'E}
893 \DeclareInputText{145}{\ae}
894 \DeclareInputText{146}{\AE}
895 \DeclareInputText{147}{\^o}
896 \DeclareInputText{148}{\"o}
897 \DeclareInputText{149}{\@tabacckludge`o}
898 \DeclareInputText{150}{\^u}
899 \DeclareInputText{151}{\@tabacckludge`u}

900 \DeclareInputText{152}{\"y}
901 \DeclareInputText{153}{\"O}
902 \DeclareInputText{154}{\"U}

903 ⟨/cp437 | cp437de | cp865⟩ 904 ⟨∗cp437 | cp437de⟩

905 \DeclareInputText{155}{\textcent}
906 \DeclareInputText{156}{\pounds}
907 \DeclareInputText{157}{\textyen}

908 ⟨/cp437 | cp437de⟩ 909 ⟨∗cp865⟩

910 \DeclareInputText{155}{\o}
911 \DeclareInputText{156}{\pounds}
912 \DeclareInputText{157}{\O}

913 ⟨/cp865⟩ 914 ⟨∗cp437 | cp437de | cp865⟩

915 \DeclareInputText{158}{\textpeseta} % Pt
916 \DeclareInputText{159}{\textflorin}
917 \DeclareInputText{160}{\@tabacckludge'a} % 160
918 \DeclareInputText{161}{\@tabacckludge'\i}
919 \DeclareInputText{162}{\@tabacckludge'o}
920 \DeclareInputText{163}{\@tabacckludge'u}
921 \DeclareInputText{164}{\~n}
922 \DeclareInputText{165}{\~N}
923 \DeclareInputText{166}{\textordfeminine}
924 \DeclareInputText{167}{\textordmasculine}
925 \DeclareInputText{168}{\textquestiondown}
926 %\DeclareInputText{169}{} % left upper corner
927 \DeclareInputMath{170}{\lnot}
928 \DeclareInputText{171}{\textonehalf}
929 \DeclareInputText{172}{\textonequarter}
930 \DeclareInputText{173}{\textexclamdown}
931 \DeclareInputText{174}{\guillemotleft}
932 ⟨cp437 | cp437de⟩\DeclareInputText{175}{\guillemotright}
933 ⟨cp865⟩\DeclareInputText{175}{\textcurrency}
934 %\DeclareInputText{176}{\textlightgraybox}
935 %\DeclareInputText{177}{\textgraybox}
936 %\DeclareInputText{178}{\textdarkgraybox}
937 %\DeclareInputMath{179}{} % vertical bar
938 %\DeclareInputText{180}{} % vertical bar with branch to left
939 %\DeclareInputText{181}{} % vertical bar with double branch to left
940 %\DeclareInputText{182}{} % double bar with single branch to left
941 %\DeclareInputText{183}{} % graphic
942 %\DeclareInputText{184}{} % graphic
943 %\DeclareInputMath{185}{} % vertical double bar with branch to left
944 %\DeclareInputMath{186}{} % vertical double bar
945 %\DeclareInputText{187}{} % double upper right corner
946 %\DeclareInputMath{188}{} % double lower right corner
947 %\DeclareInputText{189}{} % graphic
948 %\DeclareInputText{190}{} % graphic
949 %\DeclareInputMath{191}{\ensuremath{\rceil}}
950 %\DeclareInputMath{192}{\ensuremath{\lfloor}}
951 %\DeclareInputText{193}{} % dash with branch up
952 %\DeclareInputText{194}{} % dash with branch down

953 %\DeclareInputText{195}{} % vertical bar with branch to right
954 %\DeclareInputText{196}{} % horizontal bar, not endash
955 %\DeclareInputText{197}{} % vertical bar crossed with dash
956 %\DeclareInputText{198}{} % graphic
957 %\DeclareInputText{199}{} % graphic
958 %\DeclareInputText{200}{} % double lower left corner
959 %\DeclareInputText{201}{} % double upper left corner
960 %\DeclareInputText{202}{} % double dash with branch up
961 %\DeclareInputText{203}{} % double dash with branch down
962 %\DeclareInputText{204}{} % double bar with branch right
963 %\DeclareInputText{205}{=} % double dash
964 %\DeclareInputText{206}{} % double bar crossing double dash
965 %\DeclareInputText{207}{} % graphic
966 %\DeclareInputMath{208}{}
967 %\DeclareInputText{209}{}
968 %\DeclareInputText{210}{}
969 %\DeclareInputText{211}{}
970 %\DeclareInputText{212}{}
971 %\DeclareInputText{213}{}
972 %\DeclareInputText{214}{}
973 %\DeclareInputText{215}{}
974 %\DeclareInputText{216}{}
975 %\DeclareInputMath{217}{\ensuremath{\rfloor}} % lower right corner
976 %\DeclareInputMath{218}{\ensuremath{\lceil}}  % upper left corner
977 %\DeclareInputText{219}{} % black box
978 %\DeclareInputText{220}{} % lower half of black box
979 %\DeclareInputText{221}{} % left bar
980 %\DeclareInputText{222}{} % right bar
981 %\DeclareInputText{223}{} % upper half of black box
982 \DeclareInputMath{224}{\alpha}

983 ⟨/cp437 | cp437de | cp865⟩

984 ⟨cp437 | cp865⟩\DeclareInputMath{225}{\beta}
985 ⟨cp437de⟩\DeclareInputText{225}{\ss}

986 ⟨∗cp437 | cp437de | cp865⟩

987 \DeclareInputMath{226}{\Gamma}
988 \DeclareInputMath{227}{\pi}
989 \DeclareInputMath{228}{\Sigma}
990 \DeclareInputMath{229}{\sigma}
991 \DeclareInputMath{230}{\mu}
992 \DeclareInputMath{231}{\gamma}
993 \DeclareInputMath{232}{\Phi}
994 \DeclareInputMath{233}{\theta}
995 \DeclareInputMath{234}{\Omega}
996 \DeclareInputMath{235}{\delta}
997 \DeclareInputMath{236}{\infty}
998 \DeclareInputMath{237}{\phi}
999 \DeclareInputMath{238}{\varepsilon}
1000 \DeclareInputMath{239}{\cap}
1001 \DeclareInputMath{240}{\equiv}
1002 \DeclareInputMath{241}{\pm}
1003 \DeclareInputMath{242}{\ge}
1004 \DeclareInputMath{243}{\le}
1005 %\DeclareInputMath{244}{}   % upper part of integral sign
1006 %\DeclareInputMath{245}{}   % lower part of integral sign

1007 \DeclareInputMath{246}{\div}
1008 \DeclareInputMath{247}{\approx}
1009 \DeclareInputText{248}{\textdegree}
1010 \DeclareInputText{249}{\textperiodcentered}
1011 \DeclareInputText{250}{\textbullet}
1012 \DeclareInputMath{251}{\surd}
1013 \DeclareInputMath{252}{\mathnsuperior}
1014 \DeclareInputMath{253}{\mathtwosuperior}
1015 \DeclareInputText{254}{\textblacksquare} % right name?
1016 \DeclareInputText{255}{\nobreakspace}

1017 ⟨/cp437 | cp437de | cp865⟩

14 The Macintosh encodings

This input encoding was contributed by Constantin Kahn (kahn@math.toronto.edu), with minor modifications by Alan Jeffrey.

1018 ⟨∗applemac⟩

1019 \DeclareInputText{128}{\"A}
1020 \DeclareInputText{129}{\r A}
1021 \DeclareInputText{130}{\c C}
1022 \DeclareInputText{131}{\@tabacckludge'E}
1023 \DeclareInputText{132}{\~N}
1024 \DeclareInputText{133}{\"O}
1025 \DeclareInputText{134}{\"U}
1026 \DeclareInputText{135}{\@tabacckludge'a}
1027 \DeclareInputText{136}{\@tabacckludge`a}
1028 \DeclareInputText{137}{\^a}
1029 \DeclareInputText{138}{\"a}
1030 \DeclareInputText{139}{\~a}
1031 \DeclareInputText{140}{\r a}
1032 \DeclareInputText{141}{\c c}
1033 \DeclareInputText{142}{\@tabacckludge'e}
1034 \DeclareInputText{143}{\@tabacckludge`e}
1035 \DeclareInputText{144}{\^e}
1036 \DeclareInputText{145}{\"e}
1037 \DeclareInputText{146}{\@tabacckludge'\i}
1038 \DeclareInputText{147}{\@tabacckludge`\i}
1039 \DeclareInputText{148}{\^\i}
1040 \DeclareInputText{149}{\"\i}
1041 \DeclareInputText{150}{\~n}
1042 \DeclareInputText{151}{\@tabacckludge'o}
1043 \DeclareInputText{152}{\@tabacckludge`o}
1044 \DeclareInputText{153}{\^o}
1045 \DeclareInputText{154}{\"o}
1046 \DeclareInputText{155}{\~o}
1047 \DeclareInputText{156}{\@tabacckludge'u}
1048 \DeclareInputText{157}{\@tabacckludge`u}
1049 \DeclareInputText{158}{\^u}
1050 \DeclareInputText{159}{\"u}
1051 \DeclareInputText{160}{\dag}
1052 \DeclareInputText{161}{\textdegree}
1053 \DeclareInputText{162}{\textcent}

1054 \DeclareInputText{163}{\pounds}
1055 \DeclareInputText{164}{\S}
1056 \DeclareInputText{165}{\textbullet}
1057 \DeclareInputText{166}{\P}
1058 \DeclareInputText{167}{\ss}
1059 \DeclareInputText{168}{\textregistered}
1060 \DeclareInputText{169}{\copyright}
1061 \DeclareInputText{170}{\texttrademark}
1062 \DeclareInputText{171}{\@tabacckludge'{}}
1063 \DeclareInputText{172}{\"{}}
1064 \DeclareInputMath{173}{\neq}
1065 \DeclareInputText{174}{\AE}
1066 \DeclareInputText{175}{\O}
1067 \DeclareInputMath{176}{\infty}
1068 \DeclareInputMath{177}{\pm}
1069 \DeclareInputMath{178}{\leq}
1070 \DeclareInputMath{179}{\geq}
1071 \DeclareInputText{180}{\textyen}
1072 \DeclareInputMath{181}{\mu}
1073 \DeclareInputMath{182}{\partial}
1074 \DeclareInputMath{183}{\Sigma}
1075 \DeclareInputMath{184}{\Pi}
1076 \DeclareInputMath{185}{\pi}
1077 \DeclareInputMath{186}{\int}
1078 \DeclareInputText{187}{\textordfeminine}
1079 \DeclareInputText{188}{\textordmasculine}
1080 \DeclareInputMath{189}{\Omega}
1081 \DeclareInputText{190}{\ae}
1082 \DeclareInputText{191}{\o}
1083 \DeclareInputText{192}{\textquestiondown}
1084 \DeclareInputText{193}{\textexclamdown}
1085 \DeclareInputMath{194}{\lnot}
1086 \DeclareInputMath{195}{\surd}
1087 \DeclareInputText{196}{\textflorin}
1088 \DeclareInputMath{197}{\approx}
1089 \DeclareInputMath{198}{\Delta}
1090 \DeclareInputText{199}{\guillemotleft}
1091 \DeclareInputText{200}{\guillemotright}
1092 \DeclareInputText{201}{\dots}
1093 \DeclareInputText{202}{\nobreakspace}
1094 \DeclareInputText{203}{\@tabacckludge`A}
1095 \DeclareInputText{204}{\~A}
1096 \DeclareInputText{205}{\~O}
1097 \DeclareInputText{206}{\OE}
1098 \DeclareInputText{207}{\oe}
1099 \DeclareInputText{208}{\textendash}
1100 \DeclareInputText{209}{\textemdash}
1101 \DeclareInputText{210}{\textquotedblleft}
1102 \DeclareInputText{211}{\textquotedblright}
1103 \DeclareInputText{212}{\textquoteleft}
1104 \DeclareInputText{213}{\textquoteright}
1105 \DeclareInputMath{214}{\div}
1106 \DeclareInputMath{215}{\diamond}
1107 \DeclareInputText{216}{\"y}

1108 \DeclareInputText{217}{\"Y}
1109 \DeclareInputMath{218}{/}
1110 \DeclareInputText{219}{\textcurrency}
1111 \DeclareInputText{220}{\guilsinglleft}
1112 \DeclareInputText{221}{\guilsinglright}
1113 \DeclareInputText{222}{fi}
1114 \DeclareInputText{223}{fl}
1115 \DeclareInputText{224}{\ddag}
1116 \DeclareInputText{225}{\textperiodcentered}
1117 \DeclareInputText{226}{\quotesinglbase}
1118 \DeclareInputText{227}{\quotedblbase}
1119 \DeclareInputText{228}{\textperthousand}
1120 \DeclareInputText{229}{\^A}
1121 \DeclareInputText{230}{\^E}
1122 \DeclareInputText{231}{\@tabacckludge'A}
1123 \DeclareInputText{232}{\"E}
1124 \DeclareInputText{233}{\@tabacckludge`E}
1125 \DeclareInputText{234}{\@tabacckludge'I}
1126 \DeclareInputText{235}{\^I}
1127 \DeclareInputText{236}{\"I}
1128 \DeclareInputText{237}{\@tabacckludge`I}
1129 \DeclareInputText{238}{\@tabacckludge'O}
1130 \DeclareInputText{239}{\^O}
1131 \DeclareInputText{240}{\textapplelogo}
1132 \DeclareInputText{241}{\@tabacckludge`O}
1133 \DeclareInputText{242}{\@tabacckludge'U}
1134 \DeclareInputText{243}{\^U}
1135 \DeclareInputText{244}{\@tabacckludge`U}
1136 \DeclareInputText{245}{\i}
1137 \DeclareInputText{246}{\^{}}
1138 \DeclareInputText{247}{\~{}}
1139 \DeclareInputText{248}{\@tabacckludge={}}
1140 \DeclareInputText{249}{\u{}}
1141 \DeclareInputText{250}{\.{}}
1142 \DeclareInputText{251}{\r{}}
1143 \DeclareInputText{252}{\c\ }
1144 \DeclareInputText{253}{\H{}}
1145 \DeclareInputText{254}{\k\ }
1146 \DeclareInputText{255}{\v{}}

1147 ⟨/applemac⟩

This input encoding for the Apple Central European Codepage was contributed by Radek Tryc and Marcin Wolinski <wolinski@mimuw.edu.pl>.

1148 ⟨∗applemacce⟩

1149 \DeclareInputText{128}{\"A}
1150 \DeclareInputText{131}{\@tabacckludge'E}
1151 \DeclareInputText{132}{\k A}
1152 \DeclareInputText{133}{\"O}
1153 \DeclareInputText{134}{\"U}
1154 \DeclareInputText{136}{\k a}
1155 \DeclareInputText{137}{\v C}
1156 \DeclareInputText{138}{\" a}
1157 \DeclareInputText{139}{\v c}
1158 \DeclareInputText{140}{\@tabacckludge'C}

1159 \DeclareInputText{141}{\@tabacckludge'c}
1160 \DeclareInputText{143}{\@tabacckludge'Z}
1161 \DeclareInputText{144}{\@tabacckludge'z}
1162 \DeclareInputText{151}{\@tabacckludge'o}
1163 \DeclareInputText{159}{\"u}
1164 \DeclareInputText{154}{\"o}
1165 \DeclareInputText{133}{\"O}
1166 \DeclareInputText{134}{\"U}
1167 \DeclareInputText{153}{\^o}
1168 \DeclareInputText{160}{\textdagger}
1169 \DeclareInputText{161}{\textdegree}
1170 \DeclareInputText{162}{\k E}
1171 \DeclareInputText{163}{\pounds}
1172 \DeclareInputText{164}{\textsection}
1173 \DeclareInputText{165}{\textbullet}
1174 \DeclareInputText{166}{\textparagraph}
1175 \DeclareInputText{167}{\SS}
1176 \DeclareInputText{168}{\textregistered}
1177 \DeclareInputText{171}{\k e}
1178 \DeclareInputText{193}{\@tabacckludge'N}
1179 \DeclareInputText{169}{\textcopyright}
1180 \DeclareInputText{184}{\l}
1181 \DeclareInputText{196}{\@tabacckludge'n}
1182 \DeclareInputText{199}{\guillemotleft}
1183 \DeclareInputText{200}{\guillemotright}
1184 \DeclareInputText{201}{\textellipsis}
1185 \DeclareInputText{202}{\nobreakspace}
1186 \DeclareInputText{208}{\textendash}
1187 \DeclareInputText{209}{\textemdash}
1188 \DeclareInputText{210}{\textquotedbleft}
1189 \DeclareInputText{211}{\textquotedblright}
1190 \DeclareInputText{212}{\textquoteleft}
1191 \DeclareInputText{213}{\textquoteright}
1192 \DeclareInputText{214}{\textdiv}
1193 \DeclareInputText{220}{\guilsinglleft}
1194 \DeclareInputText{221}{\guilsinglright}
1195 \DeclareInputText{222}{\v r}
1196 \DeclareInputText{226}{\quotesinglbase}
1197 \DeclareInputText{227}{\quotedblbase}
1198 \DeclareInputText{229}{\@tabacckludge'S}
1199 \DeclareInputText{230}{\@tabacckludge's}
1200 \DeclareInputText{238}{\@tabacckludge'O}
1201 \DeclareInputText{239}{\^O}
1202 \DeclareInputText{251}{\.Z}
1203 \DeclareInputText{252}{\L}
1204 \DeclareInputText{253}{\.z}

1205 ⟨/applemacce⟩

15 The Next encoding

This input encoding is based on work by Stefan Ried (stef@theo-phys.uni-essen.de

  and Holger Uhr (huhr@uni-paderborn.de).

    Further extended by Jens Heise (heisbeee@calvados.zrz.TU-Berlin.DE).

1206 ⟨∗next⟩

1207 \DeclareInputText{128}{\nobreakspace}
1208 \DeclareInputText{129}{\@tabacckludge`A}
1209 \DeclareInputText{130}{\@tabacckludge'A}
1210 \DeclareInputText{131}{\^A}
1211 \DeclareInputText{132}{\~A}
1212 \DeclareInputText{133}{\"A}
1213 \DeclareInputText{134}{\r A}
1214 \DeclareInputText{135}{\c C}
1215 \DeclareInputText{136}{\@tabacckludge`E}
1216 \DeclareInputText{137}{\@tabacckludge'E}
1217 \DeclareInputText{138}{\^E}
1218 \DeclareInputText{139}{\"E}
1219 \DeclareInputText{140}{\@tabacckludge`I}
1220 \DeclareInputText{141}{\@tabacckludge'I}
1221 \DeclareInputText{142}{\^I}
1222 \DeclareInputText{143}{\"I}
1223 \DeclareInputText{144}{\DH}
1224 \DeclareInputText{145}{\~N}
1225 \DeclareInputText{146}{\@tabacckludge`O}
1226 \DeclareInputText{147}{\@tabacckludge'O}
1227 \DeclareInputText{148}{\^O}
1228 \DeclareInputText{149}{\~O}
1229 \DeclareInputText{150}{\"O}
1230 \DeclareInputText{151}{\@tabacckludge`U}
1231 \DeclareInputText{152}{\@tabacckludge'U}
1232 \DeclareInputText{153}{\^U}
1233 \DeclareInputText{154}{\"U}
1234 \DeclareInputText{155}{\@tabacckludge'Y}
1235 \DeclareInputText{156}{\TH}
1236 \DeclareInputMath{157}{\mu}
1237 \DeclareInputMath{158}{\times}
1238 \DeclareInputMath{159}{\div}
1239 \DeclareInputText{160}{\copyright}
1240 \DeclareInputText{161}{\textexclamdown}
1241 \DeclareInputText{162}{\textcent}
1242 \DeclareInputText{163}{\pounds}
1243 \DeclareInputMath{164}{/}
1244 \DeclareInputText{165}{\textyen}
1245 \DeclareInputText{166}{\textflorin}
1246 \DeclareInputText{167}{\S}
1247 \DeclareInputText{168}{\textcurrency}
1248 \DeclareInputText{169}{\textquotesingle}
1249 \DeclareInputText{170}{\textquotedblleft}
1250 \DeclareInputText{171}{\guillemotleft}
1251 \DeclareInputText{172}{\guilsinglleft}
1252 \DeclareInputText{173}{\guilsinglright}
1253 \DeclareInputText{174}{fi}
1254 \DeclareInputText{175}{fl}
1255 \DeclareInputText{176}{\textregistered}
1256 \DeclareInputText{177}{\textendash}
1257 \DeclareInputText{178}{\dag}

1258 \DeclareInputText{179}{\ddag}
1259 \DeclareInputText{180}{\textperiodcentered}
1260 \DeclareInputText{181}{\textbar}
1261 \DeclareInputText{182}{\P}
1262 \DeclareInputText{183}{\textbullet}
1263 \DeclareInputText{184}{\quotesinglbase}
1264 \DeclareInputText{185}{\quotedblbase}
1265 \DeclareInputText{186}{\textquotedblright}
1266 \DeclareInputText{187}{\guillemotright}
1267 \DeclareInputText{188}{\dots}
1268 \DeclareInputText{189}{\textperthousand}
1269 \DeclareInputMath{190}{\lnot}
1270 \DeclareInputText{191}{\textquestiondown}
1271 \DeclareInputMath{192}{\mathonesuperior}
1272 \DeclareInputText{193}{\@tabacckludge`{}}
1273 \DeclareInputText{194}{\@tabacckludge'{}}
1274 \DeclareInputText{195}{\^{}}
1275 \DeclareInputText{196}{\~{}}
1276 \DeclareInputText{197}{\@tabacckludge={}}
1277 \DeclareInputText{198}{\u{}}
1278 \DeclareInputText{199}{\.{}}
1279 \DeclareInputText{200}{\"{}}
1280 \DeclareInputMath{201}{\mathtwosuperior}
1281 \DeclareInputText{202}{\textdegree}
1282 \DeclareInputText{203}{\c\ }
1283 \DeclareInputMath{204}{\maththreesuperior}
1284 \DeclareInputText{205}{\H{}}
1285 \DeclareInputText{206}{\capitalogonek{}}
1286 \DeclareInputText{207}{\v{}}
1287 \DeclareInputText{208}{\textemdash}
1288 \DeclareInputMath{209}{\pm}
1289 \DeclareInputText{210}{\textonequarter}
1290 \DeclareInputText{211}{\textonehalf}
1291 \DeclareInputText{212}{\textthreequarters}
1292 \DeclareInputText{213}{\@tabacckludge`a}
1293 \DeclareInputText{214}{\@tabacckludge'a}
1294 \DeclareInputText{215}{\^a}
1295 \DeclareInputText{216}{\~a}
1296 \DeclareInputText{217}{\"a}
1297 \DeclareInputText{218}{\r a}
1298 \DeclareInputText{219}{\c c}
1299 \DeclareInputText{220}{\@tabacckludge`e}
1300 \DeclareInputText{221}{\@tabacckludge'e}
1301 \DeclareInputText{222}{\^e}
1302 \DeclareInputText{223}{\"e}
1303 \DeclareInputText{224}{\@tabacckludge`\i}
1304 \DeclareInputText{225}{\AE}
1305 \DeclareInputText{226}{\@tabacckludge'\i}
1306 \DeclareInputText{227}{\textordfeminine}
1307 \DeclareInputText{228}{\^\i}
1308 \DeclareInputText{229}{\"\i}
1309 \DeclareInputMath{230}{\eth}
1310 \DeclareInputText{231}{\~n}

1311 \DeclareInputText{232}{\L}
1312 \DeclareInputText{233}{\O}
1313 \DeclareInputText{234}{\OE}
1314 \DeclareInputText{235}{\textordmasculine}
1315 \DeclareInputText{236}{\@tabacckludge`o}
1316 \DeclareInputText{237}{\@tabacckludge'o}
1317 \DeclareInputText{238}{\^o}
1318 \DeclareInputText{239}{\~o}
1319 \DeclareInputText{240}{\"o}
1320 \DeclareInputText{241}{\ae}
1321 \DeclareInputText{242}{\@tabacckludge`u}
1322 \DeclareInputText{243}{\@tabacckludge'u}
1323 \DeclareInputText{244}{\^u}
1324 \DeclareInputText{245}{\i}
1325 \DeclareInputText{246}{\"u}
1326 \DeclareInputText{247}{\@tabacckludge'y}
1327 \DeclareInputText{248}{\l}
1328 \DeclareInputText{249}{\o}
1329 \DeclareInputText{250}{\oe}
1330 \DeclareInputText{251}{\ss}
1331 \DeclareInputText{252}{\th}
1332 \DeclareInputText{253}{\"y}
1333 \DeclareInputText{254}{}
1334 \DeclareInputText{255}{}

1335 ⟨/next⟩

16 The MS Windows ANSI encoding cp1252, and the ISO Latin-1 and Latin-9 encodings

The MS Windows ANSI, cp 1252, input encoding was contributed by Berthold

K.P. Horn (bkph@ai.mit.edu).

It has two very different names so the same code produces two files. These encoding files both define the characters in the MS Windows 3.1 ANSI encoding (Western Europe), also known as code page 1252, which is based on ISO Latin-1 but has important additions in the 128–159 range.

Designed for: Danish, Dutch, English, Finnish, French, German, Icelandic, Italian, Norwegian, Portuguese, Spanish, and Swedish.

Note: Windows ANSI — like Macintosh standard roman encoding — has quotesingle at 39, and grave at 96 — which is here ignored.

The ISO Latin-1 encoding file defines only the restricted range of characters available in the ISO 8859-1 encoding.

The ISO Latin-9 encoding file defines the characters in the ISO 8859-15 encoding. It was contributed by Karsten Tinnefeld (karsten@tinnefeld.com). It differs only a small amount from ISO Latin-1 and is a replacement for it that contains a few characters that are needed for French and Finnish. Further, a slot for the Euro currency sign has been added and this could be the killer argument for many 8-bit texts to be written in Latin-9 in the future.

According to a Linux manpage, ISO Latin-9 supports Albanian, Basque, Breton, Catalan, Danish, Dutch, English, Estonian, Faroese, Finnish, French, Frisian, Galician, German, Greenlandic, Icelandic, Irish Gaelic, Italian, Latin, Luxemburgish, Norwegian, Portuguese, Rhaeto-Romanic, Scottish Gaelic, Spanish and Swedish.

The characters added in latin9.def are (in L^{A}TEX notation):

       \texteuro, \v S   \v s   \v Z   \v z   \OE   \oe   \" Y

They displace the following characters from latin1.def:

         \textcurrency   \textbrokenbar   \"{}   \'{}   \c{}
         \textonequarter   \textonehalf   \textthreequarters

1336 ⟨∗cp1252⟩

1337 \DeclareInputText{128}{\texteuro}
1338 \DeclareInputText{130}{\quotesinglbase}
1339 \DeclareInputText{131}{\textflorin}
1340 \DeclareInputText{132}{\quotedblbase}
1341 \DeclareInputText{133}{\dots}
1342 \DeclareInputText{134}{\dag}
1343 \DeclareInputText{135}{\ddag}
1344 \DeclareInputText{136}{\^{}}
1345 \DeclareInputText{137}{\textperthousand}
1346 \DeclareInputText{138}{\v S}
1347 \DeclareInputText{139}{\guilsinglleft}
1348 \DeclareInputText{140}{\OE}
1349 \DeclareInputText{142}{\v Z}
1350 \DeclareInputText{145}{\textquoteleft}
1351 \DeclareInputText{146}{\textquoteright}
1352 \DeclareInputText{147}{\textquotedblleft}
1353 \DeclareInputText{148}{\textquotedblright}
1354 \DeclareInputText{149}{\textbullet}
1355 \DeclareInputText{150}{\textendash}
1356 \DeclareInputText{151}{\textemdash}
1357 \DeclareInputText{152}{\~{}}
1358 \DeclareInputText{153}{\texttrademark}
1359 \DeclareInputText{154}{\v s}
1360 \DeclareInputText{155}{\guilsinglright}
1361 \DeclareInputText{156}{\oe}
1362 \DeclareInputText{158}{\v z}
1363 \DeclareInputText{159}{\"Y}

1364 ⟨/cp1252⟩

This (somewhat confused) table is now even more disordered: first we deal with those few characters that are different in latin9, then with the rest.

1365 ⟨∗cp1252 | latin1⟩

1366 \DeclareInputText{164}{\textcurrency}
1367 \DeclareInputText{166}{\textbrokenbar}
1368 \DeclareInputText{168}{\"{}}
1369 \DeclareInputText{180}{\@tabacckludge'{}}
1370 \DeclareInputText{184}{\c\ }
1371 \DeclareInputText{188}{\textonequarter}
1372 \DeclareInputText{189}{\textonehalf}
1373 \DeclareInputText{190}{\textthreequarters}

1374 ⟨/cp1252 | latin1⟩ 1375 ⟨∗latin9⟩

1376 \DeclareInputText{164}{\texteuro}
1377 \DeclareInputText{166}{\v S}
1378 \DeclareInputText{168}{\v s}
1379 \DeclareInputText{180}{\v Z}
1380 \DeclareInputText{184}{\v z}
1381 \DeclareInputText{188}{\OE}
1382 \DeclareInputText{189}{\oe}
1383 \DeclareInputText{190}{\"Y}

1384 ⟨/latin9⟩

1385 ⟨∗cp1252 | latin1 | latin9⟩

1386 \DeclareInputText{160}{\nobreakspace}
1387 \DeclareInputText{176}{\textdegree}
1388 \DeclareInputText{161}{\textexclamdown}
1389 \DeclareInputMath{177}{\pm}
1390 \DeclareInputText{162}{\textcent}
1391 \DeclareInputMath{178}{\mathtwosuperior}
1392 \DeclareInputText{163}{\pounds}
1393 \DeclareInputMath{179}{\maththreesuperior}
1394 \DeclareInputText{165}{\textyen}
1395 \DeclareInputMath{181}{\mu}
1396 \DeclareInputText{182}{\P}
1397 \DeclareInputText{167}{\S}
1398 \DeclareInputText{183}{\textperiodcentered}
1399 \DeclareInputText{169}{\copyright}
1400 \DeclareInputMath{185}{\mathonesuperior}
1401 \DeclareInputText{170}{\textordfeminine}
1402 \DeclareInputText{186}{\textordmasculine}
1403 \DeclareInputText{171}{\guillemotleft}
1404 \DeclareInputText{187}{\guillemotright}
1405 \DeclareInputMath{172}{\lnot}
1406 \DeclareInputText{173}{\-}
1407 \DeclareInputText{174}{\textregistered}
1408 \DeclareInputText{175}{\@tabacckludge={}}
1409 \DeclareInputText{191}{\textquestiondown}
1410 \DeclareInputText{192}{\@tabacckludge`A}
1411 \DeclareInputText{208}{\DH}
1412 \DeclareInputText{193}{\@tabacckludge'A}
1413 \DeclareInputText{209}{\~N}
1414 \DeclareInputText{194}{\^A}
1415 \DeclareInputText{210}{\@tabacckludge`O}
1416 \DeclareInputText{195}{\~A}
1417 \DeclareInputText{211}{\@tabacckludge'O}
1418 \DeclareInputText{196}{\"A}
1419 \DeclareInputText{212}{\^O}
1420 \DeclareInputText{197}{\r A}
1421 \DeclareInputText{213}{\~O}
1422 \DeclareInputText{198}{\AE}
1423 \DeclareInputText{214}{\"O}
1424 \DeclareInputText{199}{\c C}
1425 \DeclareInputMath{215}{\times}
1426 \DeclareInputText{200}{\@tabacckludge`E}
1427 \DeclareInputText{216}{\O}

1428 \DeclareInputText{201}{\@tabacckludge'E}
1429 \DeclareInputText{217}{\@tabacckludge`U}
1430 \DeclareInputText{202}{\^E}
1431 \DeclareInputText{218}{\@tabacckludge'U}
1432 \DeclareInputText{203}{\"E}
1433 \DeclareInputText{219}{\^U}
1434 \DeclareInputText{204}{\@tabacckludge`I}
1435 \DeclareInputText{220}{\"U}
1436 \DeclareInputText{205}{\@tabacckludge'I}
1437 \DeclareInputText{221}{\@tabacckludge'Y}
1438 \DeclareInputText{206}{\^I}
1439 \DeclareInputText{222}{\TH}
1440 \DeclareInputText{207}{\"I}
1441 \DeclareInputText{223}{\ss}
1442 \DeclareInputText{224}{\@tabacckludge`a}
1443 \DeclareInputText{240}{\dh}
1444 \DeclareInputText{225}{\@tabacckludge'a}
1445 \DeclareInputText{241}{\~n}
1446 \DeclareInputText{226}{\^a}
1447 \DeclareInputText{242}{\@tabacckludge`o}
1448 \DeclareInputText{227}{\~a}
1449 \DeclareInputText{243}{\@tabacckludge'o}
1450 \DeclareInputText{228}{\"a}
1451 \DeclareInputText{244}{\^o}
1452 \DeclareInputText{229}{\r a}
1453 \DeclareInputText{245}{\~o}
1454 \DeclareInputText{230}{\ae}
1455 \DeclareInputText{246}{\"o}
1456 \DeclareInputText{231}{\c c}
1457 \DeclareInputMath{247}{\div}
1458 \DeclareInputText{232}{\@tabacckludge`e}
1459 \DeclareInputText{248}{\o}
1460 \DeclareInputText{233}{\@tabacckludge'e}
1461 \DeclareInputText{249}{\@tabacckludge`u}
1462 \DeclareInputText{234}{\^e}
1463 \DeclareInputText{250}{\@tabacckludge'u}
1464 \DeclareInputText{235}{\"e}
1465 \DeclareInputText{251}{\^u}
1466 \DeclareInputText{236}{\@tabacckludge`\i}
1467 \DeclareInputText{252}{\"u}
1468 \DeclareInputText{237}{\@tabacckludge'\i}
1469 \DeclareInputText{253}{\@tabacckludge'y}
1470 \DeclareInputText{238}{\^\i}
1471 \DeclareInputText{254}{\th}
1472 \DeclareInputText{239}{\"\i}
1473 \DeclareInputText{255}{\"y}

1474 ⟨/cp1252 | latin1 | latin9⟩

17 The ISO 8859-16 (Latin10) encoding

This set of coded graphic characters is intended for use in data and text processing applications and also for information interchange. The set contains graphic characters used for general purpose applications in typical office environments in at least the following languages: Albanian, Croatian, English, Finnish, French, German, Hungarian, Irish Gaelic (new orthography), Italian, Latin, Polish, Romanian, and Slovenian. This set of coded graphic characters may be regarded as a version of an 8-bit code according to ISO/IEC 2022 or ISO/IEC 4873 at level 1. [ISO 8859-16:2001(E), p. 1]

ISO 8859-16 was primarily designed for single-byte encoding the Romanian language. The UTF-8 charset is the preferred and in today’s MIME software more widely implemented encoding suitable for Romanian.

Data for the L^{A}TEX support was mainly provided by Ionel Ciobîcă with additions and corrections taken from

  http://www.unicode.org/Public/MAPPINGS/ISO8859/8859-16.TXT.

1475 ⟨∗latin10⟩

The “comma below” accent is provided here in a crude (better than nothing) version.

1476 \ProvideTextCommandDefault\textcommabelow[1]
1477 {\hmode@bgroup\ooalign{\null#1\crcr\hidewidth
1478  \raise-.31ex\hbox{\check@mathfonts

Use \sf@size instead of \ssf@size if the comma looks too small:

1479  \fontsize\ssf@size\z@
1480  \math@fontsfalse\selectfont,}\hidewidth}\egroup}
1481 \ProvideTextCommandDefault\textpm{\ensuremath\pm}
1482 \DeclareInputText{160}{\nobreakspace}
1483 \DeclareInputText{161}{\k A}
1484 \DeclareInputText{162}{\k a}
1485 \DeclareInputText{163}{\L}
1486 \DeclareInputText{164}{\texteuro}
1487 \DeclareInputText{165}{\quotedblbase}
1488 \DeclareInputText{166}{\v S}
1489 \DeclareInputText{167}{\S}
1490 \DeclareInputText{168}{\v s}
1491 \DeclareInputText{169}{\textcopyright}
1492 \DeclareInputText{170}{\textcommabelow S}
1493 \DeclareInputText{171}{\guillemotleft}
1494 \DeclareInputText{172}{\@tabacckludge'Z}
1495 \DeclareInputText{173}{\-}
1496 \DeclareInputText{174}{\@tabacckludge'z}
1497 \DeclareInputText{175}{\.Z}
1498 \DeclareInputText{176}{\textdegree}
1499 \DeclareInputText{177}{\textpm}
1500 \DeclareInputText{178}{\v C}
1501 \DeclareInputText{179}{\l}
1502 \DeclareInputText{180}{\v Z}
1503 \DeclareInputText{181}{\textquotedblright}
1504 \DeclareInputText{182}{\P}
1505 \DeclareInputText{183}{\textperiodcentered}
1506 \DeclareInputText{184}{\v z}
1507 \DeclareInputText{185}{\v c}
1508 \DeclareInputText{186}{\textcommabelow s}
1509 \DeclareInputText{187}{\guillemotright}
1510 \DeclareInputText{188}{\OE}
1511 \DeclareInputText{189}{\oe}

1512 \DeclareInputText{190}{\"Y}
1513 \DeclareInputText{191}{\.z}
1514 \DeclareInputText{192}{\@tabacckludge`A}
1515 \DeclareInputText{193}{\@tabacckludge'A}
1516 \DeclareInputText{194}{\^A}
1517 \DeclareInputText{195}{\u A}
1518 \DeclareInputText{196}{\"A}
1519 \DeclareInputText{197}{\@tabacckludge'C}
1520 \DeclareInputText{198}{\AE}
1521 \DeclareInputText{199}{\c C}
1522 \DeclareInputText{200}{\@tabacckludge`E}
1523 \DeclareInputText{201}{\@tabacckludge'E}
1524 \DeclareInputText{202}{\^E}
1525 \DeclareInputText{203}{\"E}
1526 \DeclareInputText{204}{\@tabacckludge`I}
1527 \DeclareInputText{205}{\@tabacckludge'I}
1528 \DeclareInputText{206}{\^I}
1529 \DeclareInputText{207}{\"I}
1530 \DeclareInputText{208}{\DJ}
1531 \DeclareInputText{209}{\@tabacckludge'N}
1532 \DeclareInputText{210}{\@tabacckludge`O}
1533 \DeclareInputText{211}{\@tabacckludge'O}
1534 \DeclareInputText{212}{\^O}
1535 \DeclareInputText{213}{\H O}
1536 \DeclareInputText{214}{\"O}
1537 \DeclareInputText{215}{\@tabacckludge'S}
1538 \DeclareInputText{216}{\H U}
1539 \DeclareInputText{217}{\@tabacckludge`U}
1540 \DeclareInputText{218}{\@tabacckludge'U}
1541 \DeclareInputText{219}{\^U}
1542 \DeclareInputText{220}{\"U}
1543 \DeclareInputText{221}{\k E}
1544 \DeclareInputText{222}{\textcommabelow T}
1545 \DeclareInputText{223}{\ss}
1546 \DeclareInputText{224}{\@tabacckludge`a}
1547 \DeclareInputText{225}{\@tabacckludge'a}
1548 \DeclareInputText{226}{\^a}
1549 \DeclareInputText{227}{\u a}
1550 \DeclareInputText{228}{\"a}
1551 \DeclareInputText{229}{\@tabacckludge'c}
1552 \DeclareInputText{230}{\ae}
1553 \DeclareInputText{231}{\c c}
1554 \DeclareInputText{232}{\@tabacckludge`e}
1555 \DeclareInputText{233}{\@tabacckludge'e}
1556 \DeclareInputText{234}{\^e}
1557 \DeclareInputText{235}{\"e}
1558 \DeclareInputText{236}{\@tabacckludge`\i}
1559 \DeclareInputText{237}{\@tabacckludge'\i}
1560 \DeclareInputText{238}{\^\i}
1561 \DeclareInputText{239}{\"\i}
1562 \DeclareInputText{240}{\dj}
1563 \DeclareInputText{241}{\@tabacckludge'n}
1564 \DeclareInputText{242}{\@tabacckludge`o}

1565 \DeclareInputText{243}{\@tabacckludge'o}
1566 \DeclareInputText{244}{\^o}
1567 \DeclareInputText{245}{\H o}
1568 \DeclareInputText{246}{\"o}
1569 \DeclareInputText{247}{\@tabacckludge's}
1570 \DeclareInputText{248}{\H u}
1571 \DeclareInputText{249}{\@tabacckludge`u}
1572 \DeclareInputText{250}{\@tabacckludge'u}
1573 \DeclareInputText{251}{\^u}
1574 \DeclareInputText{252}{\"u}
1575 \DeclareInputText{253}{\k e}
1576 \DeclareInputText{254}{\textcommabelow t}
1577 \DeclareInputText{255}{\"y}

1578 ⟨/latin10⟩

18 The MS Windows encoding cp1250

This is an MS Windows encoding for Central and Eastern Europe, known as Code Page 1250; it was provided by Marcin Woliński wolinski@melkor.mimuw.edu.pl and edited by Chris Rowley (see v0.94 below) to make it consistent with other files.

1579 ⟨∗cp1250⟩

1580 \DeclareInputText{128}{\texteuro}
1581 %\DeclareInputText{129}{} % n/u
1582 \DeclareInputText{130}{\quotesinglbase}
1583 %\DeclareInputText{131}{} % n/u
1584 \DeclareInputText{132}{\quotedblbase}
1585 \DeclareInputText{133}{\dots}
1586 \DeclareInputText{134}{\dag}
1587 \DeclareInputText{135}{\ddag}
1588 %\DeclareInputText{136}{} % n/u
1589 \DeclareInputText{137}{\textperthousand}
1590 \DeclareInputText{138}{\v S}
1591 \DeclareInputText{139}{\guilsinglleft}
1592 \DeclareInputText{140}{\@tabacckludge'S}
1593 \DeclareInputText{141}{\v T}
1594 \DeclareInputText{142}{\v Z}
1595 \DeclareInputText{143}{\@tabacckludge'Z}
1596 %\DeclareInputText{144}{} % n/u
1597 \DeclareInputText{145}{\textquoteleft}
1598 \DeclareInputText{146}{\textquoteright}
1599 \DeclareInputText{147}{\textquotedblleft}
1600 \DeclareInputText{148}{\textquotedblright}
1601 \DeclareInputText{149}{\textbullet}
1602 \DeclareInputText{150}{\textendash}
1603 \DeclareInputText{151}{\textemdash}
1604 %\DeclareInputText{152}{} % n/u
1605 \DeclareInputText{153}{\texttrademark}
1606 \DeclareInputText{154}{\v s}
1607 \DeclareInputText{155}{\guilsinglright}
1608 \DeclareInputText{156}{\@tabacckludge's}
1609 \DeclareInputText{157}{\v t} % t caron (t') ?

1610 \DeclareInputText{158}{\v z}
1611 \DeclareInputText{159}{\@tabacckludge'z}
1612 \DeclareInputText{160}{\nobreakspace}
1613 \DeclareInputText{161}{\v{}}
1614 \DeclareInputText{162}{\u{}}
1615 \DeclareInputText{163}{\L}
1616 \DeclareInputText{164}{\textcurrency}
1617 \DeclareInputText{165}{\k A}
1618 \DeclareInputText{166}{\textbrokenbar}
1619 \DeclareInputText{167}{\S}
1620 \DeclareInputText{168}{\"{}}
1621 \DeclareInputText{169}{\copyright}
1622 \DeclareInputText{170}{\c S}
1623 \DeclareInputText{171}{\guillemotleft}
1624 \DeclareInputMath{172}{\lnot}
1625 \DeclareInputText{173}{\-}% soft hyphen
1626 \DeclareInputText{174}{\textregistered}
1627 \DeclareInputText{175}{\.Z}
1628 \DeclareInputText{176}{\textdegree}
1629 \DeclareInputMath{177}{\pm}% plus-minus
1630 \DeclareInputText{178}{\k\ }
1631 \DeclareInputText{179}{\l}
1632 \DeclareInputText{180}{\@tabacckludge'{}}
1633 \DeclareInputMath{181}{\mu}% micro sign
1634 \DeclareInputText{182}{\P}
1635 \DeclareInputText{183}{\textperiodcentered}
1636 \DeclareInputText{184}{\c\ }
1637 \DeclareInputText{185}{\k a}
1638 \DeclareInputText{186}{\c s}
1639 \DeclareInputText{187}{\guillemotright}
1640 \DeclareInputText{188}{\v L}% L caron (L') ?
1641 \DeclareInputText{189}{\H{}}
1642 \DeclareInputText{190}{\v l}% l caron (l') ?
1643 \DeclareInputText{191}{\. z}
1644 \DeclareInputText{192}{\@tabacckludge'R}
1645 \DeclareInputText{193}{\@tabacckludge'A}
1646 \DeclareInputText{194}{\^A}
1647 \DeclareInputText{195}{\u A}
1648 \DeclareInputText{196}{\"A}
1649 \DeclareInputText{197}{\@tabacckludge'L}
1650 \DeclareInputText{198}{\@tabacckludge'C}
1651 \DeclareInputText{199}{\c C}
1652 \DeclareInputText{200}{\v C}
1653 \DeclareInputText{201}{\@tabacckludge'E}
1654 \DeclareInputText{202}{\k E}
1655 \DeclareInputText{203}{\"E}
1656 \DeclareInputText{204}{\v E}
1657 \DeclareInputText{205}{\@tabacckludge'I}
1658 \DeclareInputText{206}{\^I}
1659 \DeclareInputText{207}{\v D}
1660 \DeclareInputText{208}{\DJ} % D stroke
1661 \DeclareInputText{209}{\@tabacckludge'N}
1662 \DeclareInputText{210}{\v N}

1663 \DeclareInputText{211}{\@tabacckludge'O}
1664 \DeclareInputText{212}{\^O}
1665 \DeclareInputText{213}{\H O}
1666 \DeclareInputText{214}{\"O}
1667 \DeclareInputMath{215}{\times}
1668 \DeclareInputText{216}{\v R}
1669 \DeclareInputText{217}{\r U}
1670 \DeclareInputText{218}{\@tabacckludge'U}
1671 \DeclareInputText{219}{\H U}
1672 \DeclareInputText{220}{\" U}
1673 \DeclareInputText{221}{\@tabacckludge'Y}
1674 \DeclareInputText{222}{\c T}
1675 \DeclareInputText{223}{\ss}
1676 \DeclareInputText{224}{\@tabacckludge'r}
1677 \DeclareInputText{225}{\@tabacckludge'a}
1678 \DeclareInputText{226}{\^a}
1679 \DeclareInputText{227}{\u a}
1680 \DeclareInputText{228}{\" a}
1681 \DeclareInputText{229}{\@tabacckludge'l}
1682 \DeclareInputText{230}{\@tabacckludge'c}
1683 \DeclareInputText{231}{\c c}
1684 \DeclareInputText{232}{\v c}
1685 \DeclareInputText{233}{\@tabacckludge'e}
1686 \DeclareInputText{234}{\k e}
1687 \DeclareInputText{235}{\"e}
1688 \DeclareInputText{236}{\v e}
1689 \DeclareInputText{237}{\@tabacckludge'\i}
1690 \DeclareInputText{238}{\^\i}
1691 \DeclareInputText{239}{\v d} % d caron (d') ?
1692 \DeclareInputText{240}{\dj} % d stroke
1693 \DeclareInputText{241}{\@tabacckludge'n}
1694 \DeclareInputText{242}{\v n}
1695 \DeclareInputText{243}{\@tabacckludge'o}
1696 \DeclareInputText{244}{\^o}
1697 \DeclareInputText{245}{\H o}
1698 \DeclareInputText{246}{\"o}
1699 \DeclareInputMath{247}{\div}
1700 \DeclareInputText{248}{\v r}
1701 \DeclareInputText{249}{\r u}
1702 \DeclareInputText{250}{\@tabacckludge'u}
1703 \DeclareInputText{251}{\H u}
1704 \DeclareInputText{252}{\"u}
1705 \DeclareInputText{253}{\@tabacckludge'y}
1706 \DeclareInputText{254}{\c t}
1707 \DeclareInputText{255}{\.{}}

1708 ⟨/cp1250⟩

19 The MS Windows encoding cp1257

This is an MS Windows encoding for Baltic languages, known as Code Page 1257; it was provided on the web at http://nad.datapro.lv/LaTeX and edited by Frank to make it consistent with other files.

1709 ⟨∗cp1257⟩

1710 \DeclareInputText{160}{\nobreakspace}
1711 \DeclareInputText{161}{\k A}
1712 \DeclareInputText{163}{\c R}
1713 \DeclareInputText{164}{\textcurrency}
1714 \DeclareInputText{165}{\~I}
1715 \DeclareInputText{166}{\c L}
1716 \DeclareInputText{167}{\S}
1717 \DeclareInputText{168}{\"\ }
1718 \DeclareInputText{169}{\v S}
1719 \DeclareInputText{170}{\@tabacckludge=E}
1720 \DeclareInputText{171}{\c G}
1721 \DeclareInputText{173}{\-}
1722 \DeclareInputText{174}{\v Z}
1723 \DeclareInputText{175}{\@tabacckludge={}}
1724 \DeclareInputText{176}{\textdegree}
1725 \DeclareInputText{177}{\k a}
1726 \DeclareInputText{178}{\k\ }
1727 \DeclareInputText{179}{\c r}
1728 \DeclareInputText{180}{\@tabacckludge'{}}
1729 \DeclareInputText{181}{\~\i}
1730 \DeclareInputText{182}{\c l}
1731 \DeclareInputText{183}{\v\ }
1732 \DeclareInputText{184}{\c\ }
1733 \DeclareInputText{185}{\v s}
1734 \DeclareInputText{186}{\@tabacckludge=e}
1735 \DeclareInputText{187}{\c g}
1736 \DeclareInputText{189}{\NG}
1737 \DeclareInputText{190}{\v z}
1738 \DeclareInputText{191}{\ng}
1739 \DeclareInputText{192}{\@tabacckludge=A}
1740 \DeclareInputText{193}{\@tabacckludge'A}
1741 \DeclareInputText{194}{\=A}
1742 \DeclareInputText{195}{\~A}
1743 \DeclareInputText{196}{\"A}
1744 \DeclareInputText{197}{\r A}
1745 \DeclareInputText{198}{\AE}
1746 \DeclareInputText{199}{\=E}
1747 \DeclareInputText{200}{\v C}
1748 \DeclareInputText{201}{\@tabacckludge'E}
1749 \DeclareInputText{202}{\k E}
1750 \DeclareInputText{203}{\"E}
1751 \DeclareInputText{204}{\c G}
1752 \DeclareInputText{205}{\c K}
1753 \DeclareInputText{206}{\@tabacckludge=I}
1754 \DeclareInputText{207}{\c L}
1755 \DeclareInputText{208}{\v S}
1756 \DeclareInputText{209}{\c N}
1757 \DeclareInputText{210}{\c N}
1758 \DeclareInputText{211}{\c N}
1759 \DeclareInputText{212}{\^O}
1760 \DeclareInputText{213}{\~O}
1761 \DeclareInputText{214}{\"O}
1762 \DeclareInputMath{215}{\times}

1763 \DeclareInputText{216}{\O}
1764 \DeclareInputText{217}{\k U}
1765 \DeclareInputText{218}{\@tabacckludge'U}
1766 \DeclareInputText{219}{\@tabacckludge=U}
1767 \DeclareInputText{220}{\"U}
1768 \DeclareInputText{221}{\~U}
1769 \DeclareInputText{222}{\v Z}
1770 \DeclareInputText{223}{\ss}
1771 \DeclareInputText{224}{\@tabacckludge=a}
1772 \DeclareInputText{225}{\@tabacckludge'a}
1773 \DeclareInputText{226}{\@tabacckludge=a}
1774 \DeclareInputText{227}{\~a}
1775 \DeclareInputText{228}{\"a}
1776 \DeclareInputText{229}{\r a}
1777 \DeclareInputText{230}{\ae}
1778 \DeclareInputText{231}{\@tabacckludge=e}
1779 \DeclareInputText{232}{\v c}
1780 \DeclareInputText{233}{\@tabacckludge'e}
1781 \DeclareInputText{234}{\k e}
1782 \DeclareInputText{235}{\"e}
1783 \DeclareInputText{236}{\k g}
1784 \DeclareInputText{237}{\c k}
1785 \DeclareInputText{238}{\@tabacckludge=\i}
1786 \DeclareInputText{239}{\k l}
1787 \DeclareInputText{240}{\v s}
1788 \DeclareInputText{241}{\c n}
1789 \DeclareInputText{242}{\k n}
1790 \DeclareInputText{243}{\c k}
1791 \DeclareInputText{244}{\^o}
1792 \DeclareInputText{245}{\~o}
1793 \DeclareInputText{246}{\"o}
1794 \DeclareInputMath{247}{\div}
1795 \DeclareInputText{248}{\o}
1796 \DeclareInputText{249}{\k u}
1797 \DeclareInputText{250}{\@tabacckludge'u}
1798 \DeclareInputText{251}{\@tabacckludge=u}
1799 \DeclareInputText{252}{\"u}
1800 \DeclareInputText{253}{\~u}
1801 \DeclareInputText{254}{\v z}
1802 \DeclareInputText{255}{\.\ }

1803 ⟨/cp1257⟩

Finally, we reset the category code of the at sign at the end of all .def files.

1804 ⟨−package⟩\makeatother